### v1

# Differential forms: Lagrange interpolation, sampling and approximation on polynomial admissible integral $k$-meshes

Abstract.
In this work we address the problem of interpolating and approximating differential forms starting from data defined by integration. We show that many aspects of nodal interpolation can naturally be carried to this more general framework; in contrast, some of them require the introduction of geometric and measure theoretic hypotheses. After characterizing the norms of the operators involved, we introduce the concept of admissible integral $k$-mesh, which allows for the construction of robust approximation schemes, and is used to extract interpolation sets with high stability properties. To this end, the concepts of Fekete currents and Leja sequences of currents are formalized, and a numerical scheme for their approximation is proposed.
1. Introduction
Differential forms are generalizations of functions endowed with rich geometric and algebraic structure. An evocative definition of differential forms is due to Flanders, who describe them as “things which occur under the integral sign” in [32, p. 1]. This very informal definition stresses how deeply the development of differential forms is intertwined with integration theory, and it should hence not sound surprising that they are in fact a main character of geometric measure theory.
A motivated use of differential forms as a mathematical tool can be traced, from the middle of the nineteenth century, in the work of Poincaré, Pfaff, Jacobi, and Deahna and was justified by earlier findings of, among others, Gauss, Clairault, Euler, Stokes and Lagrange, see [40]. Such early studies on differential forms were mainly motivated by the mathematical modelization of physical phenomena like electromagnetism, fluid dynamics or continuum mechanics. Those pioneering works had in fact a great impact on physicists and, at the same time, drew relevant inspiration from physics: for instance, one may see Gauss’ law for the electric field as a special case of Stokes’ Theorem. The rigorous foundation and a modern systematic mathematical treatment of the theory of differential forms was developed only in the twentieth century by the seminal work of Cartan [25], de Rham [28], and Hodge [37], mostly with a geometrical perspective. Later, differential forms spread from the realm of differential geometry to other branches of mathematics, becoming a standard tool in many topics such as PDEs, complex variables, and geometric measure theory.
Nowadays, differential forms are commonly presented as sections of the $k$-th exterior power of a cotangent bundle. Hence, they represent vector valued “functions” endowed with an alternating product. This very rigid algebraic structure is what makes differential forms tailored for integration and clarifies why they relate the local and the global behavior of physical phenomena. As a consequence, differential forms are generally exploited to embody the geometry of the domain under consideration into the physics of problems: this perspective is influencing the numerical community as well. In fact, differential forms are currently being employed in finite elements approximation [4], mimetic methods [43] and, more generally, in structure preserving methods for PDEs [36]. The extensive and profitable use in such contexts lead mostly to a study of convergence in Sobolev norms (see, e.g., [33]), which is suggested by regularity theory and functional analysis. On the contrary, a thorough study of continuous differential forms under the light of integration and uniform approximation theory is still lacking. The principal aim of this paper is to sneak into and partially fill such a gap.
First of all, we remark that the projection of differential forms onto a finite dimensional subspace is in fact the common ground to all the aforementioned numerical methods. When interpolation is enforced by weights [58], one in essence obtains a generalization of Lagrange interpolation where nodal evaluations are replaced by certain currents of integration [31, p. 381]. These linear functionals are in fact represented by an oriented geometrical support in the physical space where integration of the differential form is performed; they hence assume the meaning of circulations along lines (e.g., for the electric field), fluxes across faces (e.g., for the Stokes’ flow or the magnetic field), and so on. Likewise, other projection methods such as least squares fitting [62, 19] involve the minimization of suitable norms based on integration; note that the physical interpretation does not change.
In order to study the approximation of differential forms from a quantitative perspective, a concept of distance is needed. Remarkably, it is possible to introduce a norm on the space of differential $k$-forms with continuous coefficients $\mathscr{D}_{0}^{k}$ which on the one hand hinges on integration, and on the other plays the role that uniform norm has for functions. This norm, denoted by $\|\cdot\|_{0}$ and discussed in Section 2.1, turns the space of differential $k$-forms with continuous coefficients into a Banach space.
While the possibility of approximating a differential form $\omega$ hinges on the regularity of its coefficients, in practice the quality of the actual approximation obtained by a linear projection $L$ onto a finite dimensional subspace $\mathscr{V}$ depends also on the norm of the operator $L$. One in particular has the Lebesgue type estimate of the error $\|\omega-L\omega\|_{0}\leq\|L\|_{\mathrm{op}}\inf\{\|\omega-\eta\|_{0},\,\eta\in\mathscr{V}\}.$ As one leaves the nodal framework, this norm $\|L\|_{\mathrm{op}}$ becomes difficult to compute and even to characterize. Hence, in the case of interpolation one estimates such a norm from above by the generalized Lebesgue constant $\mathscr{L}$, first introduced in [3]. The same situation is faced in least squares fitting, where the norm of the discrete least squares projection operator can generally only be estimated from above by the quantity $\mathscr{M}$ defined in Eq. (23), which is the natural counterpart of $\mathscr{L}$. This issue is treated in Section 3, where we exhibit a measure theoretical condition that grants the equalities $\|\Pi\|_{\mathrm{op}}=\mathscr{L}$, for the interpolation operator $\Pi$, and $\|P\|_{\mathrm{op}}=\mathscr{M}$ for the discrete least squares projector $P$.
To continue the parallelism with the nodal setting, we observe that both the quantities $\mathscr{L}$ and $\mathscr{M}$ relate the approximation scheme with the aforementioned $k$-dimensional oriented supports of integration. When considering a diffeomorphism $\varphi$ of the $n$-dimensional reference domain, the $k$-dimensional measure of the supports of integration does not simply scale with the determinant of $D\varphi$. Drawing inspiration from finite elements estimates [29, §11.2], we consider the compositions $\Pi\circ\varphi^{*}$ and $P\circ\varphi^{*}$ of the projection operators with the pullback $\varphi^{*}$ induced by $\varphi$. Working in such a context, in Sections 3.1.1 and 3.2.2 we investigate the dependence of $\|\Pi\circ\varphi^{*}\|_{\mathrm{op}}$ and $\|P\circ\varphi^{*}\|_{\mathrm{op}}$ on the singular values of $D\varphi$. The results we obtain, reported in Theorem 2 and Proposition 2, extend those of [1] in three directions: the order $k$ of the form, the shape of the considered supports, and the consideration of general diffeomorphisms instead of invertible affine mappings.
In the context of uniform polynomial approximation of functions, classical problems such as the selection of “good” interpolation points, the construction of low cardinality discrete least squares projectors, positive quadrature, and polynomial optimization, have been very profitably tackled by coupling a sampling inequality with a bound on the cardinality of the sampling set [12, 53, 55]. This approach, first proposed in [24], lead to the introduction of the concept of admissible mesh, that is, a sequence of uniform norming sets whose cardinality grows polynomially with respect to the considered degree. The aim of Section 4 is to extend this idea to the framework of differential $k$-forms with polynomial coefficients $\mathscr{P}\Lambda^{k}$. In Definition 2 we thus introduce the concept of admissible integral $k$-meshes. Such a generalization extends the nodal case under several points of view: first, the uniform norm is replaced by $\|\cdot\|_{0}$, and further, integral currents supported on pieces of $k$-dimensional affine varieties are considered in place of nodal evaluations. The construction of admissible integral $k$-meshes is addressed in the remaining part of Section 4. Exploiting the Markov Inequality first, and the Baran Inequality then, we provide two explicit strategies for the construction of admissible integral $k$-meshes. The first option is very flexible in the sense that the presented construction can be applied to quite general reference domains, but presents a high cardinality of the obtained mesh as main drawback. On the contrary, the second technique provides low cardinality meshes, but can only be applied to the few reference domains for which the Baran metric is explicitly known.
With the concept of admissible integral $k$-meshes at hand one can extend the procedures developed in the nodal context. In the case of interpolation, one can extract Fekete currents [22] and Leja sequences from currents supported on the admissible mesh instead of generic currents supported on the reference domain. This approach yields extremely robust interpolation schemes, as the norm of the corresponding projection operator is a priori controlled by the dimension of the range of the projection. This construction is detailed in Section 5.1 and gives the first example in the literature of interpolation schemes for polynomial differential forms whose generalized Lebesgue constants grows at most at a polynomial rate with respect to the polynomial degree. Further, approximate Fekete and Leja currents may be computed via numerical linear algebra factorizations [13, 12]. Despite the heuristic nature of this approach, numerical tests for the nodal case showed its high effectiveness.
In the case of discrete least squares fitting, admissible integral $k$-meshes also yield an estimate of the norm of the operator $P$ projecting onto polynomial forms $\mathscr{P}_{r}\Lambda^{k}$. In Proposition 5 it is in fact exhibited that $\|P\|_{\mathrm{op}}$ has polynomial growth with respect to the polynomial degree $r$. Remarkably, a stronger result can be obtained mimicking the construction carried out in [24].
2. Background and tools
2.1. Test forms, norms, integration, and currents of order zero
Let us denote denote by $\Lambda_{k}$ the $k$-th exterior power of $\mathbb{R}^{n}$. This vector space is constructed by taking $k$ copies of $\mathbb{R}^{n}$ and considering the quotient $\mathbb{R}^{n}\otimes\dots\otimes\mathbb{R}^{n}/I$, where $I$ is the ideal generated by repeating terms $v\otimes v$. On this space there is a natural alternating structure that relates its elements via the wedge product: if $\sigma$ is a permutation on $k$ objects, then
$\Lambda_{k}\ni v:=v_{1}\wedge\ldots\wedge v_{k}=\mathrm{sgn}(\sigma)\,v_{\sigma(1)}\wedge\ldots\wedge v_{\sigma(k)}.$
By construction, $\Lambda_{k}$ is the $\binom{n}{k}$-dimensional vector space spanned by wedge products of the form $e_{\alpha}:=e_{\alpha(1)}\wedge\ldots\wedge e_{\alpha(k)}$, where $e_{i}$ is the $i$-th element of the cardinal basis of $\mathbb{R}^{n}$ and $\alpha$ is any increasing mapping $\{1,\ldots,k\}\to\{1,\ldots,n\}$. We consider the lexicographical on such a basis. Consistently, any element $v$ of $\Lambda_{k}$ can be written as
$v={\sum_{|\alpha|=k}}^{\prime}\,a_{\alpha}e_{\alpha},$
where $a_{\alpha}$ is a real coefficient and the prime denotes summation over increasing multi-indices of length $k$. The dual space of $\Lambda_{k}$ is $\Lambda^{k}:=\left(\Lambda_{k}\right)^{*}$, and represents the $k$-th exterior power of the space of linear forms on $\mathbb{R}^{n}$. The dual pairing $\langle\omega_{1}\wedge\dots\wedge\omega_{k};\,v_{1}\wedge\dots\wedge v_{k}\rangle=\det([\omega_{i}(v_{j})]_{i,j=1,\dots,k})$ between $\Lambda^{k}$ and $\Lambda_{k}$ is inherited from the symmetric (or Euclidean) tensor product $(e_{\alpha},e_{\beta})=\delta_{\alpha,\beta}$, see, e.g., [10, p. 16]. Notice that this induces a concept of Euclidean norm on the spaces $\Lambda_{k}$ and $\Lambda^{k}$.
A differential $k$-form $\omega$ is a mapping that associates to each $x\in\mathbb{R}^{n}$ a $k$-covector of $\Lambda^{k}$. Let $E\subset\mathbb{R}^{n}$ be the closure of a bounded Lipschitz domain. For any $k\in\{0,1,\dots,n\}$ we denote by $\mathscr{D}_{0}^{k}(E):=\left(\mathscr{C}^{0}(E,{\Lambda}^{k}),\|\cdot\|_{0}\right)$ the Banach space of bounded test forms of order $k$ with continuous coefficients over $E$ endowed by the norm
[…]
where $\mathcal{H}^{k}$ denotes the $k$-dimensional Hausdorff measure. Note that writing $\mathcal{H}^{k}(S)$ we are implicitly requiring that $S$ is indeed $\mathcal{H}^{k}$-measurable and has finite $\mathcal{H}^{k}$ measure.
We recall that a set $S\subset\mathbb{R}^{n}$ is said to be $k$-rectifiable if it can be written as the union of a set having zero $k$-dimensional Hausdorff measure and a countable union of Lipschitz images of $\mathbb{R}^{k}$ [42, Def. 5.4.1]. When $S$ is $\mathcal{H}^{k}$-measurable and $k$-rectifiable, then it is possible [42, Prop. 5.4.3] to write $S=S_{\mathrm{reg}}\cup S_{\mathrm{sing}}$, where the regular part $S_{\mathrm{reg}}$ of $S$ is a countable disjoint union of (measurable pieces of) $\mathscr{C}^{1}$ $k$-dimensional embedded submanifolds of $\mathbb{R}^{n}$, $\mathcal{H}^{k}(S_{\mathrm{sing}})=0$, and $S_{\mathrm{reg}}\cap S_{\mathrm{sing}}=\emptyset$.
An orientation for the $\mathcal{H}^{k}$-measurable $k$-rectifiable set $S$ is the datum of a $k$-vector field $S\ni x\mapsto\tau(x)\in\Lambda_{k}$ such that its restriction to $S_{\mathrm{reg}}$ is continuous and, for any $x\in S_{\mathrm{reg}}$, $\tau(x)=\tau_{1}(x)\wedge\dots\wedge\tau_{k}(x)$ is a Euclidean unit simple $k$-vector and $\{\tau_{1}(x),\ldots,\tau_{k}(x)\}$ is a basis of the tangent space $T_{x}S_{\mathrm{reg}}$ of $S_{\mathrm{reg}}$ at $x$. The $\mathcal{H}^{k}$-measurable $k$-rectifiable set $S$ is said to be orientable if we can find an orientation for it, and oriented if it is endowed by an orientation.
Finally, we define the integration of the continuous form $\omega\in\mathscr{C}^{0}(E,\Lambda^{k})$ over the $\mathcal{H}^{k}$-measurable, $k$-rectifiable set $S\subset E$ with respect to the orientation $\tau$ of $S$ by setting
$\int_{S}\omega:=\int_{S}\langle\omega(x);\tau(x)\rangle\mathrm{d}\mathcal{H}^{k}(x)\,.$
Remark 1.
Integration of differential forms on oriented rectifiable sets is a key feature our construction. In fact, this approach bridges the gap between the tools customarily used in interpolation and the ones of geometric measure theory. The definitions of integration and orientation we give here are standard in the context of geometric measure theory, but readers with a major background in differential geometry might find them not completely sound. Indeed, while the above definitions arise as generalizations of their differential geometry counterparts, there are few subtle differences when moving from one context to the other. We provide in Appendix A an account of the derivation of the above definitions from the smooth manifold setting.
Despite the integration appearing in Eq. (1), the norm $\|\cdot\|_{0}$ should be regarded as an appropriate generalization of the uniform norm to the space of differential forms, as the following example shows.
Example 1.
When $k=0$, $\mathcal{H}^{k}$ is the counting measure and any oriented $k$-rectifiable set $S$ is a finite collection of points $\{\xi_{i}\}_{i=1}^{M}$ in $E$ with orientation $\tau_{i}=\pm 1$ for $i=1,\ldots,M$. Further, since $\Lambda^{0}=\mathbb{R}$, $\omega\in\mathscr{D}_{0}^{0}(E)$ is a real-valued continuous function. One has
$\|\omega\|_{0}=\sup\left\{\frac{1}{M}\left|\sum_{i=1}^{M}\tau_{i}\omega(\xi_{i})\right|,\ M\in\mathbb{N},\ \xi_{i}\in E,\ \tau_{i}=\pm 1\right\}\,=\sup_{\xi\in E}\left|\omega(\xi)\right|=\|\omega\|_{\infty}.$
When $k=n$, $\mathcal{H}^{k}$ coincides with the $n$-dimensional Lebesgue measure, and $\Lambda_{k}=\operatorname*{span}{\{e_{1}\wedge\ldots\wedge e_{n}\}}$. As a consequence, one represents $\omega\in\mathscr{D}_{0}^{n}(E)$ as $\omega=f\mathrm{d}x_{1}\wedge\ldots\wedge\mathrm{d}x_{n}$ for some $f\in\mathscr{C}^{0}(E)$. We recall that the regular part $S_{\mathrm{reg}}$ of an oriented $n$-rectifiable set $S$ can be written as the countable disjoint union of connected oriented open sets $S_{i}$; this follows by combining [42, Def. 5.4.1] with [42, Prop. 5.4.3]. By the mean value theorem applied to each $S_{i}$, there exists $\xi_{i}\in S_{i}$ such that
$\frac{1}{\mathcal{H}^{n}(S)}\left|\int_{S}\omega\right|=\frac{1}{\mathcal{H}^{n}(S)}\left|\sum_{i=1}^{+\infty}\pm\int_{S_{i}}\omega\right|=\left|\sum_{i=1}^{+\infty}\pm\frac{\mathcal{H}^{n}(S_{i})}{\mathcal{H}^{n}(S)}f(\xi_{i})\right|.$
The latter term coincides with the uniform norm of $f$ over $S$, since the closure of any set $S_{i}$ is oriented and $k$-rectifiable. Thus
$\|\omega\|_{0}=\sup_{\xi\in E}\left|f(\xi)\right|=\|f\|_{\infty}.$
Working by simplicial approximation, it is not difficult to prove that one can replace the family of all rectifiable oriented sets in the definition of Eq. (1) with the collection
$\mathcal{S}^{k}(E):=\Big{\{}\text{all }k\text{-dimensional oriented simplices lying in }E\Big{\}}\,.$
Indeed, the inequality $\|\omega\|_{0}\geq\sup_{S\in\mathcal{S}^{k}(E)}(\mathcal{H}^{k}(S))^{-1}\left|\int_{S}\omega\right|$ follows directly from definition of Eq. (1), and, shrinking simplices to points, one gets
$\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\omega\right|\geq\max_{x\in E}|\omega(x)|^{*}\,,$
where
$|\omega(x)|^{*}:=\sup\{|\langle\omega(x);\tau\rangle|,\;\tau\in\Lambda_{k}\text{ simple },|\tau|\leq 1\}$
is termed comass norm of the $k$-covector $\omega(x)$. Notice that such a norm emerges naturally in this context, since a differential form is pointwise paired in (1) only with simple vectors. On the other hand, by the Hölder Inequality, one proves that $\|\omega\|_{0}\leq\max_{x\in E}|\omega(x)|^{*}$, whence
(2) $\|\omega\|_{0}=\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\omega\right|=\max_{x\in E}|\omega(x)|^{*}.$
Remark 2.
It worth saying here that the identity (2) depends on the regularity assumption on $E$. The hypothesis of $E$ being the closure of a bounded Lipschitz domain is in particular a sufficient condition and rather natural for our construction, however far from being necessary. Indeed, it is only necessary that for any $\tau\in\Lambda_{k}$ with $|\tau|=1$, $x\in E$, and $\varepsilon>0$, we can find $x_{\varepsilon}\in E$ and $\tau^{\varepsilon}:=\tau_{1}^{\varepsilon}\wedge\dots\wedge\tau_{k}^{\varepsilon}\in\Lambda_{k}$ such that $|x-x_{\varepsilon}|<\varepsilon,$ $|\tau-\tau^{\varepsilon}|<\varepsilon$, and $\{x_{\varepsilon}+\sum_{i=1}^{k}\tau_{i}^{\varepsilon}y_{i},\,y\in\varepsilon\Sigma_{k}\}\subset E.$
Example 2.
Let $E=\mathbb{S}^{1}\subset\mathbb{R}^{2}$. Consider the differential $1$-form $\omega:E\rightarrow\Lambda^{1}$ defined by $\omega:=x_{1}\mathrm{d}x_{1}+x_{2}\mathrm{d}x_{2}$, and denote by $\iota^{*}$ the pullback of the inclusion map $\iota:E\hookrightarrow\mathbb{R}^{2}$. By passing to polar coordinates, one readily sees that $\iota^{*}\omega=0$, thus
$\int_{S}\omega=\int_{S}\iota^{*}\omega=0\quad\forall S\subset\mathbb{S}^{1},$
or equivalently, $\langle\omega(x);\tau\rangle=0$ for any $x\in E$ and any vector $\tau$ tangent to $S_{\mathrm{reg}}$ at $x$. Hence, $\omega$ is the zero element of the Banach space $\mathscr{D}^{1}_{0}(E).$ In contrast, one has
$|\omega(x)|^{*}=1,\quad\forall x\in E,$
since the supremum defining the comass norm of the covector $\omega(x)$ is taken among all $1$-vectors of $\mathbb{R}^{2}$, which is a larger space than the tangent space of $\mathbb{S}^{1}$ at $x$.
Conversely, if $E=\{x\in\mathbb{R}^{2}:|x|\leq 1\},$ then $\|\omega\|_{0}=\max_{x\in E}|\omega(x)|^{*}=1.$ Indeed, for any $\varepsilon>0$ we may consider $S_{\varepsilon}=\{(1-\varepsilon+t,0)\in E:t\in[0,\varepsilon]\}$, and notice that
$\|\omega\|_{0}\geq\sup_{\varepsilon>0}\frac{1}{\mathcal{H}^{1}(S_{\varepsilon})}\left|\int_{S_{\varepsilon}}\omega\right|=\sup_{\varepsilon>0}\frac{1-(1-\varepsilon)^{2}}{2\varepsilon}=\sup_{\varepsilon>0}\left(1-\frac{\varepsilon}{2}\right)=1=\max_{x\in E}|\omega(x)|^{*}.$
The topological dual $\mathscr{M}_{0}^{k}(E):=(\mathscr{D}_{0}^{k}(E))^{\prime}$ of $\mathscr{D}_{0}^{k}(E)$ is the space of currents of order zero and dimension $k$. This space is naturally endowed by the operator norm
(3) $M(T):=\sup\{|T(\omega)|,\;\omega\in\mathscr{D}_{0}^{k}(E),\;\|\omega\|_{0}\leq 1\},$
which is referred to as the mass of the current $T$. We will denote by $\mathscr{S}_{0}^{k}(E)$ the unit sphere of $\mathscr{D}_{0}^{k}(E)$ with respect to the norm (3), i.e.
(4) $\mathscr{S}_{0}^{k}(E):=\{\omega\in\mathscr{D}_{0}^{k}(E),\,M(\omega)=1\}\,.$
If $T\in\mathscr{M}_{0}^{k}(E)$, the support $\operatorname*{supp}T$ of the current $T$ is defined as the complement in $E$ of the set where $T$ vanishes identically, i.e.,
$\operatorname*{supp}T:=\left(\bigcup\{A:\,A\text{ is open, }T(\omega)=0\,\forall\omega\in\mathscr{D}_{0}^{k}(A)\}\right)^{c}.$
In what follows we will consider in $\mathscr{M}_{0}^{k}(E)$ the subclass $\mathcal{I}^{k}(E)$ of currents of integration, i.e., any current $[S]$ defined by integration over an oriented rectifiable set $S$
(5) $[S](\omega)=\int_{S}\omega\,,$
and the subclass $\mathscr{A}^{k}(E)$ of currents of integral averaging, i.e. integral averages of the form
(6) $T_{S}(\omega):=\frac{[S](\omega)}{\mathcal{H}^{k}(S)}=\frac{1}{\mathcal{H}^{k}(S)}\int_{S}\omega\,,$
where $S$ is any $\mathcal{H}^{k}$-measurable $k$-rectifiable oriented subset of $E$. Note that in particular one has $M([S])=\mathcal{H}^{k}(S)$, and $M(T_{S})=1$ by construction. Let us remark that, unwinding the definitions of $\mathscr{A}^{k}(E)$ and $\|\cdot\|_{0}$, we get a characterization of $\|\omega\|_{0}$ in terms of averaging currents:
$\|\omega\|_{0}=\sup_{[S]\in\mathcal{I}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\omega\right|=\sup_{T\in\mathscr{A}^{k}(E)}T(\omega),\quad\forall\omega\in\mathscr{D}_{0}^{k}(E).$
Remark 3 (Averaging currents are not pullback invariant).
Recall that any diffeomorphism $\varphi:E\to\widehat{E}:=\varphi(E)$ induces a pullback $\varphi^{*}:\mathscr{D}_{0}^{k}(\widehat{E})\to\mathscr{D}_{0}^{k}(E)$ such that
(7) $\int_{\varphi(S)}\omega=\int_{S}\varphi^{*}\omega\quad\forall\omega\in\mathscr{D}_{0}^{k}(\widehat{E}),\;\forall S\subseteq E,$
provided $S$ is orientable and $k$-rectifiable. The support $\varphi(S)$ inherits an orientation from $S$ via $\varphi$. In particular, we can define (see, e.g., [42, §7.4.2]) the pushforward of currents of integration as
$\varphi_{*}[S](\omega):=[\varphi(S)](\omega)=\int_{\varphi(S)}\omega=\int_{S}\varphi^{*}\omega=[S](\varphi^{*}\omega)\,$
so $\varphi_{*}[S]\in\mathcal{I}^{k}(E)$. In contrast, applying this to (6), one finds
$\varphi_{*}T_{S}(\omega)=T_{S}(\varphi^{*}\omega)=\frac{[S](\varphi^{*}\omega)}{\mathcal{H}^{k}(S)}=\frac{[\varphi S](\omega)}{\mathcal{H}^{k}(S)}=\frac{\mathcal{H}^{k}(\varphi(S))}{\mathcal{H}^{k}(S)}T_{\varphi(S)}(\omega)\,.$
Hence $\varphi_{*}T_{S}\notin\mathscr{A}^{k}(E).$
2.2. The interpolation problem and the fitting problem
Let $\mathscr{V}(E)$ be a $N$-dimensional linear subspace of $\mathscr{D}_{0}^{k}(E)$. A set $\mathcal{T}:=\{T_{1},\dots,T_{M}\}\subset\mathscr{M}_{0}^{k}(E)$ is termed determining for $\mathscr{V}(E)$ whenever the condition $T_{i}(\omega)=0$, $i=1,2,\dots,M$ implies $\omega=0$; note that necessarily we have $M\geq N$ in such a case.
If $\mathcal{T}$ is a $\mathscr{V}(E)$-determinig set of currents, we may define an interpolation operator $\Pi:\mathscr{D}_{0}^{k}(E)\to\mathscr{V}(E)$ by asking that
(8) $T_{i}(\omega)=T_{i}(\Pi\omega),\quad i=1,\ldots,M.$
The interpolation problem (8) is indeed well-posed under the additional assumption that the dimension of the space $\operatorname*{span}\{T_{1},\dots,T_{M}\}$ is not larger than $N$, hence in particular when $M=N$. In the latter case, the set $\mathcal{T}$ is said to be unisolvent for $\mathscr{V}(E)$.
Unisolvence depends both on the space $\mathscr{V}(E)$ and the set $\mathcal{T}$, compare e.g. [4] with [18] or [23]. However, we will use the terminology unisolvent also refered to the set of supports, when the considered currents are currents of integration or normalized currents of integration as in (5) and (6).
Given any set of determining currents $\mathcal{T}:=\{T_{1},\dots,T_{M}\}$ for $\mathscr{V}(E)$ and a positive set of weights $\{w_{1},\dots,w_{M}\}$ we introduce the scalar product $(\cdot\,,\cdot)_{\mathcal{T},w}$ on $\mathscr{V}(E)$ by setting
(9) $(\omega,\eta)_{\mathcal{T},w}:=\sum_{i=1}^{M}w_{i}T_{i}(\omega)T_{i}(\theta)\,.$
Note that $(\omega,\omega)_{\mathcal{T},w}=\sum_{i=1}^{M}(\sqrt{w_{i}}T_{i}(\omega))^{2}=0$ implies $T_{i}(\omega)=0$ for all $i=1,\dots,M$ and thus $\omega=0$ since $\mathcal{T}$ is determining. Notice also that Eq. (9) still defines a non-negative simmetric bilinear form on $\mathscr{D}_{0}^{k}(E)$. The weighted discrete least squares projector $P:\mathscr{D}_{0}^{k}(E)\rightarrow\mathscr{V}(E)$,
(10) $P\omega:=\operatorname*{argmin}\{\|\omega-\theta\|_{\mathcal{T},w}^{2}:=(\omega-\theta,\omega-\theta)_{\mathcal{T},w},\;\theta\in\mathscr{V}(E)\},\quad\forall\omega\in\mathscr{D}_{0}^{k}(E),$
is then well-defined due to Pythagorean Theorem.
Considering any basis $\mathcal{V}=\{v_{j}\}_{j=1}^{N}$ of $\mathscr{V}(E)$ and a set of currents $\mathcal{T}=\{T_{i}\}_{i=1}^{N}$ we can form the Vandermonde matrix
$\operatorname*{vdm}(\mathcal{T},\mathcal{V})_{i,j}:=T_{i}(v_{j})\,,\quad i=1,\dots,M,\;j=1,\dots,N\,.$
Obviously the standard relations of Vandermonde matrices with the interpolation and the fitting problems still hold in this generalized setting. Indeed, $\mathcal{T}$ is determinig for $\mathscr{V}(E)$ if and only if $\operatorname*{vdm}(\mathcal{T},\mathcal{V})$ has full-column rank (for any basis $\mathcal{V}$), while $\mathcal{T}=\{T_{1},\dots,T_{N}\}$ is unisolvent if and only if $\operatorname*{vdm}(\mathcal{T},\mathcal{V})$ is invertible (in particular $\det\operatorname*{vdm}(\mathcal{T},\mathcal{V})\neq 0$). In such a case $\Pi\omega$ can be conveniently represented in the Lagrange form
(11) $\Pi\omega=\sum_{i=1}^{N}T_{i}(\omega)\omega_{i}\,,\quad T_{i}(\omega_{j})=\delta_{i,j}\,.$
In this case, we call $\{\omega_{1},\ldots,\omega_{N}\}$ the Lagrange basis relative to $\{T_{1},\dots,T_{N}\}$ and satisfies the explicit relationship $\omega_{i}=\sum_{j=1}^{N}[\operatorname*{vdm}(\mathcal{T},\mathcal{V})]^{-1}_{j,i}v_{j}.$
Further, we can compute an orthonormal basis $\{\eta_{1},\dots,\eta_{N}\}$ of $\mathscr{V}(E)$ (with respect to the product defined in (9)) using, e.g., Gram-Schmidt ortogonalization of $\sqrt{w_{i}}\operatorname*{vdm}(\mathcal{T},\mathcal{V})_{i,j}$, $i=1,\dots,M$ and $j=1,\dots,N.$ Then we can write
(12) $P\omega=\sum_{h=1}^{N}(\omega,\eta_{h})_{\mathcal{T},w}\eta_{h}=\sum_{h=1}^{N}\sum_{i=1}^{M}w_{i}T_{i}(\omega)T_{i}(\eta_{h})\eta_{h}\,,\quad(\eta_{i},\eta_{j})_{\mathcal{T},w}=\delta_{i,j}.$
It is clear that the opertor $\Pi$ can be regarded as a particular case of $P$, in which the weights are $w_{i}\equiv 1$, and $M=N$.
2.3. Spaces of polynomial differential forms
In this section we introduce two important instances of the finite dimensional space $\mathscr{V}(E)\subset\mathscr{D}_{0}^{k}(E)$: the space of complete polynomial forms and that of trimmed polynomial differential forms. The results of Sections 4 and 5 will be specialized to these settings.
We perform the construction of $\mathscr{V}$ and its basis on $\mathbb{R}^{n}$, and define $\mathscr{V}(E)$ as the restriction of $\mathscr{V}$ to $E$. Let us denote by $\mathscr{P}_{r}$ and $\mathscr{H}_{r}$ the space of $n$-variate polynomials of total degree $r$ and the space of homogeneous polynomial of degree $r$, respectively. Let us also define the ring of polynomials as $\mathscr{P}:=\oplus_{r=0}^{\infty}\mathscr{H}_{r}$. On each $\mathscr{H}_{r}$, we may consider the lexicographically ordered monomial basis $\mathcal{B}^{\mathrm{hom}}_{r}$. This choice induces on the monomials $x^{\beta}:=\prod_{i=1}^{n}x_{i}^{\beta_{i}}$ the graded lexicographical order. We denote by $\mathcal{B}^{\mathrm{mon}}$ such a basis for $\mathscr{P}$, and by $\mathcal{B}^{\mathrm{mon}}_{r}$ the basis for $\mathscr{P}_{r}$ consisting of the first $\binom{n+r}{r}$ elements of $\mathcal{B}^{\mathrm{mon}}$.
The space of complete polynomial differential forms is
(13) $\mathscr{P}\Lambda^{k}:=\bigoplus_{j=0}^{\infty}\mathscr{H}_{j}\otimes\Lambda^{k}=\mathscr{P}\otimes\Lambda^{k},$
that is, polynomial sections of the $k$-th exterior power of the cotangent bundle of $\mathbb{R}^{n}$. Truncating the above direct sum to $j=r$, we obtain the spaces $\mathscr{P}_{r}\Lambda^{k}$, whose dimension is readily computed as
$N(r):=\dim\mathscr{P}_{r}\Lambda^{k}=\binom{r+n}{r}\binom{n}{k}.$
Notice that, under the assumption we made on $E$, the space $\mathscr{P}_{r}\Lambda^{k}(E)$ is an $N(r)$-dimensional subspace of $\mathscr{D}_{0}^{k}{(E)}$. Since $\Lambda^{k}=\operatorname*{span}\{\mathrm{d}x_{\alpha},\,|\alpha|=k,\,\alpha\text{ increasing}\}$, for each $j$ we may define a basis $\mathcal{B}^{\mathrm{hom},k}_{j}:=\{h\otimes\mathrm{d}x_{\alpha},\,h\in\mathcal{B}^{\mathrm{hom}}_{j},\,|\alpha|=k,\alpha\text{ increasing}\}$ for $\mathscr{H}_{j}\otimes\Lambda^{k}$, where we again consider the lexicographical order on the pair $(h,\mathrm{d}x_{\alpha})$. Note also that $\{\mathcal{B}^{\mathrm{hom},k}_{j}\}_{j\in\mathbb{N}}$ induces a graded basis $\mathcal{B}^{\mathrm{mon},k}$ for $\mathscr{P}\Lambda^{k}$ via (13). Finally, truncating this basis at $N(r)$ we obtain the basis $\mathcal{B}^{\mathrm{mon},k}_{r}$ for $\mathscr{P}_{r}\Lambda^{k}$.
Example 3.
In the case $n=2$, $k=1$, and $r=1$, the above sets are:
$\mathcal{B}^{\mathrm{hom}}_{0}=\{1\},\;$ $\mathcal{B}^{\mathrm{hom},1}_{0}=\{\mathrm{d}x_{1},\mathrm{d}x_{2}\},$
$\mathcal{B}^{\mathrm{hom}}_{1}=\{x_{1},x_{2}\},\;$ $\mathcal{B}^{\mathrm{hom},1}_{1}=\{x_{1}\mathrm{d}x_{1},x_{1}\mathrm{d}x_{2},x_{2}\mathrm{d}x_{1},x_{2}\mathrm{d}x_{2}\},$
$\mathcal{B}^{\mathrm{hom}}_{2}=\{x_{1}^{2},x_{1}x_{2},x_{2}^{2}\},\;$ $\mathcal{B}^{\mathrm{hom},1}_{2}=\{x_{1}^{2}\mathrm{d}x_{1},x_{1}^{2}\mathrm{d}x_{2},x_{1}x_{2}\mathrm{d}x_{1},x_{1}x_{2}\mathrm{d}x_{2},x_{2}^{2}\mathrm{d}x_{1},x_{2}^{2}\mathrm{d}x_{2}\}\,,$
so that
$\mathcal{B}^{\mathrm{mon},1}_{2}=\{\mathrm{d}x_{1},\mathrm{d}x_{2},x_{1}\mathrm{d}x_{1},x_{1}\mathrm{d}x_{2},x_{2}\mathrm{d}x_{1},x_{2}\mathrm{d}x_{2},x_{1}^{2}\mathrm{d}x_{1},x_{1}^{2}\mathrm{d}x_{2},x_{1}x_{2}\mathrm{d}x_{1},x_{1}x_{2}\mathrm{d}x_{2},x_{2}^{2}\mathrm{d}x_{1},x_{2}^{2}\mathrm{d}x_{2}\}.$
As a second relevant space that fits our framework, we consider the space of trimmed polynomial differential forms. This space is defined as
$\mathscr{P}_{r}^{-}\Lambda^{k}:=\mathscr{P}_{r-1}\Lambda^{k}\oplus\kappa\left(\mathscr{H}_{r-1}\Lambda^{k+1}\right),$
where $\kappa:\mathscr{D}_{0}^{k}\to\mathscr{D}_{0}^{k-1}$ is the Koszul differential [46, p. 852], i.e., the contraction with the identity vector field. With respect to Cartesian coordinates, if $\omega:=p(x_{1},\ldots,x_{n})\mathrm{d}x_{\sigma(1)}\wedge\ldots\wedge\mathrm{d}x_{\sigma(k)}$, then
$\kappa\omega=\sum_{i=1}^{k}(-1)^{i}p(x_{1},\ldots,x_{n})x_{\sigma(i)}\mathrm{d}x_{\sigma(1)}\wedge\ldots\wedge{\mathrm{d}x}_{\sigma(i-1)}\wedge{\mathrm{d}x}_{\sigma(i+1)}\wedge\ldots\wedge\mathrm{d}x_{\sigma(k)}.$
In the lowest order, i.e. for $r=1$, $\mathscr{P}_{1}^{-}\Lambda^{k}(E)$ is the space of Whitney forms [63, p. 139]. For $r>1$, $\mathscr{P}_{r}^{-}\Lambda^{k}$ is an intermediate space between $\mathscr{P}_{r-1}\Lambda^{k}$ and $\mathscr{P}_{r}\Lambda^{k}$. As a consequence, most of our results cast for the complete space directly apply to the trimmed space as well.
Definition 1 (Lower triangular basis).
Let $\mathcal{Q}=\{q_{1},q_{2},\dots\}$ be a basis for $\mathscr{P}\Lambda^{k}$. We say that $\mathcal{Q}$ is lower triangular if for any $N\in\mathbb{N}$ there exist a lower triangular matrix $L^{(N)}$ such that
$q_{i}=\sum_{j=1}^{N}L^{(N)}_{i,j}b_{j}=\sum_{j=1}^{i}L^{(N)}_{i,j}b_{j}\,,$
where $\mathcal{B}^{\mathrm{mon},k}=\{b_{1},b_{2},\dots\}.$
Example 4 (Orthonormality of Whitney forms).
Let $E\subset\mathbb{R}^{n}$ be an $n$-simplex spanned by vertices $\{v_{0},\ldots,v_{n}\}$. Let $\{\lambda_{0},\ldots,\lambda_{n}\}$ be the corresponding barycentric coordinates. Let $\alpha$ be a increasing multiindex of length $k$, the Whitney $k$-form associated with a $k$-face $E_{\alpha}$ of $E$ is
$\omega_{\alpha}:=\sum_{i=0}^{k}(-1)^{i}\lambda_{\alpha(i)}\mathrm{d}\lambda_{\alpha(0)}\wedge\ldots\wedge{\mathrm{d}\lambda}_{\alpha(i-1)}\wedge{\mathrm{d}\lambda}_{\alpha(i+1)}\wedge\ldots\wedge\mathrm{d}\lambda_{\alpha(k)}.$
One may prove that
$\mathscr{P}_{1}^{-}\Lambda^{k}(E)=\operatorname*{span}\left\{\omega_{\alpha},\,E_{\alpha}\text{ is a $k$-face of $E$}\right\}$
and, letting $T_{\theta}$ denote the averaging current with respect to $E_{\theta}$,
$[E_{\theta}](\omega_{\alpha})=\mathcal{H}^{k}(E_{\theta})T_{E_{\theta}}(\omega_{\alpha})=\int_{E_{\theta}}\omega_{\alpha}=\mathcal{H}^{k}(E_{\theta})\delta_{\alpha,\theta}.$
The basis $\left\{\omega_{\alpha},\,E_{\alpha}\text{ is a $k$-face of $E$}\right\}$ of Whitney $k$-forms is orthonormal with respect to the scalar product (9) induced by the $k$-faces $E_{\alpha}$ of $E$ with weights $w_{\theta}=1$ for each $\theta$. Indeed:
$(\omega_{\alpha},\omega_{\tau})_{\mathcal{T},w}=\sum_{\theta=1}^{M}w_{\theta}T_{\theta}(\omega_{\alpha})T_{\theta}(\omega_{\tau})=\delta_{\alpha,\tau}.$
3. Operator norm and standard error estimates
Let $L:\mathscr{D}_{0}^{k}(E)\to\mathscr{V}(E)$ be a linear projector. The operator norm $\|L\|_{\mathrm{op}}\doteq\sup_{\|\omega\|_{0}\neq 0}\frac{\|L\omega\|_{0}}{\|\omega\|_{0}}=\sup_{\|\omega\|_{0}=1}\|L\omega\|_{0}$ controls the stability of the approximation procedure, in the sense that
$\|L\omega-L\widetilde{\omega}\|_{0}=\|L(\omega-\widetilde{\omega})\|_{0}\leq\|L\|_{\mathrm{op}}\|\omega-\widetilde{\omega}\|_{0}.$
Further, the Lebesgue inequality holds for each $\eta\in\mathscr{V}(E)$:
$\|L\omega-\omega\|_{0}=\|L\omega-L\eta+L\eta-\omega\|_{0}=\|L\omega-\eta+\eta-\omega\|_{0}=\|L(\omega-\eta)-(\omega-\eta)\|_{0}\leq(1+\|L\|_{\mathrm{op}})\|\omega-\eta\|_{0},$
and taking the minimum of the right hand side, we obtain
$\|\omega-L\omega\|_{0}\leq(1+\|L\|_{\mathrm{op}})\min_{\eta\in\mathscr{V}(E)}\|\omega-\eta\|_{0}.$
Note that one may expand $\omega=\sum^{\prime}\omega_{\alpha}(x)\mathrm{d}x_{\alpha}$, $\eta=\sum^{\prime}\eta_{\alpha}(x)\mathrm{d}x_{\alpha}$ and, using Eq. (2), one has
$\|\omega-\eta\|_{0}=\max_{x\in E}|\omega(x)-\eta(x)|^{*}\leq\max_{x\in E}\left({\sum_{|\alpha|=k}}^{\prime}(\omega_{\alpha}(x)-\eta_{\alpha}(x))^{2}\right)^{1/2}\leq\sqrt{\binom{n}{k}}\max_{|\alpha|=k}\|\omega_{\alpha}-\eta_{\alpha}\|_{E},$
where $\|\cdot\|_{E}$ denotes the uniform norm over $E$. Hence
$\|\omega-\eta\|_{0}\leq\min_{\eta\in\mathscr{V}(E)}\sqrt{\binom{n}{k}}\max_{|\alpha|=k}\|\omega_{\alpha}-\eta_{\alpha}\|_{E}.$
Also, when the space $\mathscr{V}(E)$ is built by tensor product $\mathscr{V}(E)=V(E)\otimes\Lambda^{k}$, then it is possible to swap the $\min$ and $\max$ operators in the estimate above to obtain
$\|\omega-\eta\|_{0}\leq\sqrt{\binom{n}{k}}\max_{|\alpha|=k}\min_{\eta_{\alpha}\in V(E)}\|\omega_{\alpha}-\eta_{\alpha}\|_{E}.$
Finally, if $V$ is a polynomial space, as in the case of $\mathscr{V}=\mathscr{P}_{r}\Lambda^{k}$ described in Section 2.3, then the term $\|\omega_{\alpha}-\eta_{\alpha}\|_{E}$ can be estimated by means of, e.g., Jackson-type theorems (further asking $\omega_{\alpha}\in\mathscr{C}^{s}(E)$, see [56]), or Bernstein-Walsh-Siciak results (for $\omega_{\alpha}$ locally holomorphic near $E$, see [47]).
Without any assumption on $\mathscr{V}(E)$, in Section 3.1 we characterize the above operator norm in the interpolation case $L=\Pi$, while the case of the fitting operator $L=P$ is studied in Section 3.2.
3.1. Interpolation: the Lebesgue constant
Suppose to be given the $\mathscr{V}(E)$-unisolvent set $\mathcal{T}:=\{T_{1},\ldots,T_{N}\}\subset\mathscr{A}^{k}(E)$ and let $S_{i}:=\operatorname*{supp}(T_{i})$. Let $\{\omega_{1},\ldots,\omega_{N}\}$ be the corresponding Lagrange basis defined in Eq. (11). Then we define
(14) $\mathscr{L}\left(\mathcal{T},\mathscr{V}(E)\right):=\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{N}|T(\omega_{i})|$
as the Lebesgue constant of the problem. In what follows we will suppress the depence on $\mathcal{T}$ and $\mathscr{V}(E)$ when clear from the context.
Remark 4.
The quantity (14) may be manipulated as follows:
$\mathscr{L}$ $=\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{N}|T(\omega_{i})|=\sup_{S\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}\left|T_{S}(\omega_{i})\right|=\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\sum_{i=1}^{N}\left|[S](\omega_{i})\right|$
$=\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\sum_{i=1}^{N}\mathcal{H}^{k}(S_{i})\left|[S]\left(\frac{\omega_{i}}{\mathcal{H}^{k}(S_{i})}\right)\right|=\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\sum_{i=1}^{N}\mathcal{H}^{k}(S_{i})\left|\int_{S}\eta_{i}\right|,$
where $\eta_{i}:=\omega_{i}/\mathcal{H}^{k}(S_{i})$ is the $i$-th element of the Lagrange basis associated with the currents of integration $[S_{i}]$, see [21]. In the simplicial framework, the latter quantity of the above chain of equalities was already termed Lebesgue constant in [3].
We show how $\mathscr{L}$ is related with the norm $\|\Pi\|_{\mathrm{op}}$.
Proposition 1.
Let $\Pi:\mathscr{D}_{0}^{k}(E)\rightarrow\mathscr{V}(E)$ be the interpolation operator (11) associated with the $\mathscr{V}(E)$-unisolvent set $\mathcal{T}:=\{T_{1},\dots,T_{N}\}$. One has
$\|\Pi\|_{\mathrm{op}}\leq\mathscr{L}\,.$
Proof.
Let us denote by $\{\omega_{1},\dots,\omega_{N}\}$ the Lagrange basis of $\mathscr{V}(E)$ relative to $\mathcal{T}$. Expanding definitions given in Eq. (11) and Eq. (14), and using the triangle inequality, one computes
$\|\Pi\|_{\mathrm{op}}$ $=\sup_{\|\omega\|_{0}=1}\|\Pi\omega\|_{0}=\sup_{\|\omega\|_{0}=1}\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\Pi\omega\right|=\sup_{\|\omega\|_{0}=1}\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\sum_{i=1}^{N}T_{i}(\omega)\omega_{i}\right|$
$\leq\sup_{\|\omega\|_{0}=1}\sup_{S\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}|T_{i}(\omega)|\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\omega_{i}\right|=\sup_{\|\omega\|_{0}=1}\sup_{S\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}|T_{i}(\omega)|\left|T_{S}(\omega_{i})\right|.$
Since $\|\omega\|_{0}=1$ and $T_{i}\in\mathscr{A}^{k}$, one has that $|T_{i}(\omega)|\leq 1$ for each $i$. Then $\|\Pi\|_{\mathrm{op}}\leq\mathscr{L}$. ∎
The distance between $\|\Pi\|_{\mathrm{op}}$ and $\mathscr{L}$ has been observed, in the case $T_{i}=T_{S_{i}}$, when the family of the supports $S_{i}$’s presents a significant overlapping, see [20, Fig. 3]. In the opposite scenario, i.e. when there is some distance between the supports, one may exploit bump forms [17, p. 25] and construct $\omega\in\mathscr{D}_{0}^{k}(E)$ with $\|\omega\|_{0}=1$ for which equality in the above proof is attained. The intermediate case $\mathcal{H}^{k}(S_{i}\cap S_{j})=0$ for all $i$ and $j$ is more involved, but we are still able to prove the following.
Theorem 1.
Let $\mathcal{T}=\{T_{1},\dots,T_{N}\}\subset\mathscr{A}^{k}(E)$ is a $\mathscr{V}(E)$-unisolvent set such that $\mathcal{H}^{k}(\operatorname*{supp}T_{i}\cap\operatorname*{supp}T_{j})=0$ for $i\neq j$. Then the operator $\Pi:\mathscr{D}_{0}^{k}(E)\rightarrow\mathscr{V}(E)$ satisfies
$\|\Pi\|_{\mathrm{op}}=\mathscr{L}.$
Proof.
Due to Proposition 1, it suffices to prove that, under the additional hypothesis $\mathcal{H}^{k}(\operatorname*{supp}T_{i}\cap\operatorname*{supp}T_{j})=0$ for $i\neq j$, the inequality $\|\Pi\|_{\mathrm{op}}\geq\mathscr{L}$ holds. To do so, we show that for each $0<\varepsilon<\mathscr{L}$ there exists $\omega_{\varepsilon}\in\mathscr{D}_{0}^{k}(E)$ such that
$\|\Pi\omega_{\varepsilon}\|_{0}\geq\mathscr{L}-\varepsilon.$
Let us define $S_{i}:=\operatorname*{supp}T_{i}$, $\mathcal{S}:=\bigcup_{i=1}^{N}S_{i}$, and $\mathcal{N}:=\bigcap_{i=1}^{N}S_{i}$. Observe that $\mathscr{H}^{k}(\mathcal{N})=0$. For each $\varepsilon>0$, we consider a collection of relatively open subsets $U_{i}^{\varepsilon}\subset(S_{i})_{\mathrm{reg}}\setminus\mathcal{N}$ such that $\mathcal{H}^{k}(U_{i}^{\varepsilon})\geq\left(1-\frac{\varepsilon}{\mathscr{L}}\right)\mathcal{H}^{k}(S_{i})$ and an open set $\mathcal{U}\subset E\setminus\mathcal{S}$, then define the subset $\Lambda^{k}_{\varepsilon}(E)\subset\mathscr{D}_{0}^{k}(E)$ setting
$\Lambda^{k}_{\varepsilon}(E)\doteq\left\{\omega\in\mathscr{D}_{0}^{k}(E)\ :\ \omega\bigl{|}_{U_{i}^{\varepsilon}}=\pm\mathrm{d}{\operatorname*{Vol}}_{U_{i}^{\varepsilon}},\ \omega\bigl{|}_{\mathcal{U}}=0,\,\operatorname*{sgn}(\langle\omega;\tau^{i}\rangle)\text{ is constant in }S_{i},\ \|\omega\|_{0}\leq 1\right\},$
where $\tau^{i}$ is the orientation of $S_{i}.$ By definition, for each $\omega\in\Lambda^{k}_{\varepsilon}(E)$, one has
(15) $T_{S_{i}}(\omega)=\frac{1}{\mathcal{H}^{k}(S_{i})}\left|\int_{S_{i}}\omega\right|\geq\frac{1}{\mathcal{H}^{k}(S_{i})}\left|\int_{U_{i}^{\varepsilon}}\omega\right|=\frac{\mathcal{H}^{k}(U_{i}^{\varepsilon})}{\mathcal{H}^{k}(S_{i})}\geq 1-\frac{\varepsilon}{\mathscr{L}}.$
Also, $\|\omega\|_{0}=1$ for each $\omega\in\Lambda^{k}_{\varepsilon}(E)$: as $U_{i}^{\varepsilon}$ is open (in particular $\mathcal{H}^{k}(U_{i}^{\varepsilon})>0$) in $S_{i}\setminus\mathcal{S}$, we can pick a $k$-rectifiable relatively compact subset $\bar{s}\subset U_{i}^{\varepsilon}$, $\mathcal{H}^{k}(\bar{s})>0$, such that
$T_{\bar{s}}(\omega)=\frac{1}{\mathcal{H}^{k}(\bar{s})}\int_{\bar{s}}\omega=\frac{1}{\mathcal{H}^{k}(\bar{s})}\int_{\bar{s}}\pm\mathrm{d}{\operatorname*{Vol}}_{U_{i}^{\varepsilon}}=\pm\frac{1}{\mathcal{H}^{k}(\bar{s})}\int_{\bar{s}}\mathrm{d}{\operatorname*{Vol}}_{U_{i}^{\varepsilon}}=\pm 1,$
whence $\|\omega\|_{0}=1$. One thus has
$\|\Pi\|_{\mathrm{op}}\geq\sup_{\omega\in\Lambda^{k}_{\varepsilon}(E)}\|\Pi\omega\|_{0}\,.$
We now expand $\|\Pi\omega\|_{0}$ with respect to the Lagrange basis $\{\omega_{1},\dots,\omega_{N}\}$:
$\|\Pi\omega\|_{0}$ $=\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\left(\sum_{i=1}^{N}T_{i}(\omega)\omega_{i}\right)\right|=\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\sum_{i=1}^{N}T_{i}(\omega)\int_{S}\omega_{i}\right|$
$=\sup_{S\in\mathcal{S}^{k}(E)}\left|\sum_{i=1}^{N}T_{i}(\omega)T_{S}(\omega_{i})\right|=\sup_{S\in\mathcal{S}^{k}(E)}\left|\sum_{i=1}^{N}T_{S_{i}}(\omega)\ \mathrm{sgn}\ T_{S}(\omega_{i})\left|T_{S}(\omega_{i})\right|\right|.$
By the definition of $\Lambda^{k}_{\varepsilon}(E)$, for each $S\in\mathcal{S}^{k}(E)$ there exists $\omega_{\varepsilon}\in\Lambda^{k}_{\varepsilon}(S)$ such that $\mathrm{sgn}\ T_{S}(\omega_{i})T_{i}(\omega_{\varepsilon})\geq 0$ for each $S_{i}\in\mathcal{S}$. For such an $\omega_{\varepsilon}$, one then has $\mathrm{sgn}\ T_{S}(\omega_{i})T_{i}(\omega_{\varepsilon})=\left|T_{i}(\omega_{\varepsilon})\right|$ whence, by applying the lower bound (15) and some manipulation,
$\|\Pi\omega_{\varepsilon}\|_{0}$ $=\sup_{S\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}\left|T_{i}(\omega_{\varepsilon})\right|\left|T_{S}(\omega_{i})\right|\geq\sup_{S\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}\left(1-\frac{\varepsilon}{\mathscr{L}}\right)\left|T_{S}(\omega_{i})\right|$
$=\left(1-\frac{\varepsilon}{\mathscr{L}}\right)\sup_{S\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}\left|T_{S}(\omega_{i})\right|=\left(1-\frac{\varepsilon}{\mathscr{L}}\right)\mathscr{L}=\mathscr{L}-\varepsilon.$
Since $\|\Pi\|_{\mathrm{op}}\geq\|\Pi\omega_{\varepsilon}\|_{0}\geq\mathscr{L}-\varepsilon$, this proves the claim. ∎
Remark 5.
Since $\mathcal{H}^{0}(x)=1$ when $x$ is a point, Theorem 1 extends the well known characterization of the nodal Lebesgue constant $\mathscr{L}=\sup_{x}\sum_{i=1}^{N}|\ell_{i}(x)|$, these being the cardinal functions $\ell_{i}(x_{j})=\delta_{i,j}$.
3.1.1. Dependence of $\mathscr{L}$ on the reference domain
Most results regarding Lagrange interpolation of differential forms, usually applied to polynomial differential forms, are cast in terms of the standard $n$-simplex, see e.g. [18]. While this is not restrictive in terms of unisolvence (see the forthcoming Lemma 1 for a precise statement), it affects the computation of the Lebesgue constant $\mathscr{L}$, as noted in [2, §6], since the terms in Eq. (14) do not scale accordingly to each other under the action of a non isometric map.
Example 5.
Let $\widehat{E}\subseteq\mathbb{R}^{n}$ be a $k$-simplex spanned by vertices $\{v_{0},\ldots,v_{k}\}$; with a slight abuse of notation, we still denote by $\widehat{E}$ the matrix whose columns are the vectors $\{v_{0},\ldots,v_{k}\}$. Then $\mathcal{H}^{k}(\widehat{E})$, which is in fact the $k$-volume of $\widehat{E}$, is
$\mathcal{H}^{k}(\widehat{E})=\frac{1}{k!}\sqrt{\det(\widehat{E}^{\top}\widehat{E})}.$
Let $\varphi:\widehat{E}\to E:=\varphi(\widehat{E})$ be an affine mapping with invertible linear part $A\in\mathrm{GL}(n,\mathbb{R})$. Denoting by $E$ the matrix whose columns are the vectors $\{\varphi(v_{0}),\ldots,\varphi(v_{k})\}$, one has
$\mathcal{H}^{k}(E)=\frac{1}{k!}\sqrt{\det(E^{\top}E)}=\frac{1}{k!}\sqrt{\det((A\widehat{E})^{\top}A\widehat{E})}=\frac{1}{k!}\sqrt{\det(\widehat{E}^{\top}A^{\top}A\widehat{E})}.$
Denoting by $\sigma_{1}\geq\ldots\geq\sigma_{n}$ the singular values of the matrix $A$ and performing an SVD decomposition (see [34, Thm. 2.4.1] or [61, p. 854]) of such a matrix, one finds that
$\left(\prod_{i=n-k+1}^{n}\sigma_{i}\right)\mathcal{H}^{k}(\widehat{E})\leq\mathcal{H}^{k}(E)\leq\left(\prod_{i=1}^{k}\sigma_{i}\right)\mathcal{H}^{k}(\widehat{E}),$
which in general yields the equality only when $k=0$ or $k=n$.
Via the Area Formula (see, e.g., [30, §3.3]), Example 5 can be extended to any compact $\widehat{E}$, provided that $\varphi$ is a $\mathscr{C}^{1}$ diffeomorphism. Indeed, let $\iota_{E}:E\hookrightarrow D$ and $\iota_{\widehat{E}}:\widehat{E}\hookrightarrow D$ denote the inclusion maps. This gives a commuting diagram
[…]
which induces another commuting diagram
${\mathscr{V}(\mathbb{R}^{n})}$${\mathscr{V}(\widehat{E})}$${\mathscr{V}(E)\,.}$$\scriptstyle{\iota^{*}_{E}}$$\scriptstyle{\iota^{*}_{\widehat{E}}}$$\scriptstyle{\varphi^{*}}$
Since $\varphi$ is a diffeomorphism, if the set $\{\omega_{1},\ldots,\omega_{N}\}$ is a basis for $\mathscr{V}(E)$, then $\{\varphi^{*}\omega_{1},\ldots,\varphi^{*}\omega_{N}\}$ spans an $N$-dimensional vector space $\mathscr{V}(\widehat{E})$. The map $\varphi$ preserves unisolvence as its pullback $\varphi^{*}$ preserves cardinal bases associated with non-averaged currents described in Eq. (5).
Lemma 1.
The set $\{\omega_{1},\ldots,\omega_{N}\}$ is the cardinal basis for $\mathscr{V}(E)$ associated with $\{\varphi(S_{1}),\ldots,\varphi(S_{N})\}$ if and only if the set $\{\varphi^{*}\omega_{1},\ldots,\varphi^{*}\omega_{N}\}$ is the cardinal basis for $\mathscr{V}(\widehat{E})$ associated with $\{S_{1},\ldots,S_{N}\}$.
Proof.
This is immediate from the change of variable formula (7), computing
$\delta_{i,j}=\int_{\varphi(S_{i})}\omega_{j}=\int_{S_{i}}\varphi^{*}\omega_{j}.$
The claim follows from the uniqueness of the Lagrange basis. ∎
Remark 6.
Despite Lemma 1 does not hold anymore in the case of averaged currents introduced in Eq. (6), unisolvence is preserved by images of diffeomorphisms. Further, working as in Remark 3, one immediately sees that if $\omega_{i}$ is the cardinal function associated with the averaged current $T_{S_{i}}$, then $\frac{\mathcal{H}^{k}(S_{i})}{\mathcal{H}^{k}(\varphi(S_{i}))}\varphi^{*}(\omega_{i})$ is the cardinal function associated with the averaged current $T_{\varphi(S_{i})}$.
Lemma 2.
Let $S\subset\mathbb{R}^{n}$ be a $k$-rectifiable set such that $\mathcal{H}^{k}(S)<+\infty$. Let $U$ be a neighbourhood of $S$ in $\mathbb{R}^{n}$ and let $\varphi:U\rightarrow\mathbb{R}^{n}$ be a $\mathscr{C}^{1}$ diffeomorphism on its image. Then one has
(16) $\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}\mathcal{H}^{k}(S)\leq\mathcal{H}^{k}(\varphi(S))\leq\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}\mathcal{H}^{k}(S)\,,$
where $\sigma_{j}^{(\varphi)}:U\rightarrow(0,+\infty)$ are the singular values functions of the matrix $D\varphi(\cdot)$ arranged in non-increasing order, i.e., for any $x\in U$, one has $\sigma_{l}^{(\varphi)}(x)\geq\sigma_{j}^{(\varphi)}(x)$ for $l>j$.
Proof.
Recall that $S=S_{\mathrm{sing}}\bigcup\left(\cup_{j=1}^{\infty}K_{j}\right)$, where $S_{\mathrm{sing}}$ has zero $\mathcal{H}^{k}$-measure, $K_{i}\cap K_{j}=\emptyset$ for distinct indices, and the $K_{i}$’s are measurable pieces of the $\mathscr{C}^{1}$ $k$-submanifolds $S_{i}\subset\mathbb{R}^{n}$. For each $S_{i}$ consider an atlas $\{(\phi_{\alpha}^{i}:U_{\alpha}^{i}\rightarrow V_{\alpha}^{i})\}$, $U_{\alpha}^{i}\subset S_{i}$, $V_{\alpha}^{i}\subset\mathbb{R}^{k}$, and a partition of unity $\{\rho_{\alpha}^{i}\}$ subordinated to the atlas, i.e., $\rho_{\alpha}^{i}\in\mathscr{C}^{\infty}_{c}(U_{\alpha}^{i},[0,1])$, $\sum_{\alpha}\rho_{\alpha}^{i}=1$ on $S_{i}$. Finally, let $\psi_{\alpha}^{i}:=(\phi_{\alpha}^{i})^{-1}:V_{\alpha}^{i}\rightarrow U_{\alpha}^{i}$.
Since $\varphi$ is a $\mathscr{C}^{1}$ diffeomorphism, it induces an atlas of $\varphi(S_{i})$ and the partition of unity $\tilde{\rho}_{\alpha}^{i}:=\rho_{\alpha}^{i}\circ\varphi^{-1}$. We can compute $\mathcal{H}^{k}(\varphi(K_{i}))$ by the Area Formula:
$\mathcal{H}^{k}(\varphi(K_{i}))=$ $\int_{\varphi(K_{i})}\mathrm{d}\mathcal{H}^{k}=\sum_{\alpha}\int_{\varphi(U_{\alpha}^{i}\cap K_{i})}\tilde{\rho}_{\alpha}^{i}(\tilde{x})\mathrm{d}\mathcal{H}^{k}(\tilde{x})$
$=$ $\sum_{\alpha}\int_{V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\rho_{\alpha}^{i}(\psi_{\alpha}^{i}(y))\llbracket D(\varphi\circ\psi_{\alpha}^{i})\rrbracket(y)\mathrm{d}\mathcal{H}^{k}(y)$
$=$ $\sum_{\alpha}\int_{V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\rho_{\alpha}^{i}(\psi_{\alpha}^{i}(y))\llbracket D\varphi(\psi_{\alpha}^{i}(y))D\psi_{\alpha}^{i}(y)\rrbracket\mathrm{d}\mathcal{H}^{k}(y)$
$=$ $\sum_{\alpha}\int_{V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\rho_{\alpha}^{i}(\psi_{\alpha}^{i}(y))\sqrt{\det\left[(D\psi_{\alpha}^{i}(y))^{\top}(D\varphi(\psi_{\alpha}^{i}(y)))^{\top}D\varphi(\psi_{\alpha}^{i}(y))D\psi_{\alpha}^{i}(y)\right]}\mathrm{d}\mathcal{H}^{k}(y)\,.$
Consider the singular value decomposition $D\psi_{\alpha}^{i}(y)=U(y)\Sigma(y)V(y)^{\top}$. Let
$\Sigma(y)^{\top}=[\Sigma_{0}^{\top}(y)|\mathbb{O}_{k,n-k}]:=[\mathrm{diag}(\sigma_{1}(y),\dots\sigma_{k}(y))|\mathbb{O}_{k,n-k}]\,,$
where $\mathbb{O}_{k,n-k}$ is the zero matrix of size $k\times(n-k)$. Then we compute
$\sqrt{\det\left[(D\psi_{\alpha}^{i}(y))^{\top}(D\varphi(\psi_{\alpha}^{i}(y)))^{\top}D\varphi(\psi_{\alpha}^{i}(y))D\psi_{\alpha}^{i}(y)\right]}$
$=$ $\sqrt{\det(V(y)\Sigma^{\top}(y)U^{\top}(y)(D\varphi(\psi_{\alpha}^{i}(y)))^{\top}D\varphi(\psi_{\alpha}^{i}(y))U(y)\Sigma(y)V(y)^{\top})}$
$=$ $\sqrt{\det(\Sigma_{0}^{\top}(y)\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k}\Sigma_{0}(y))}$
$=$ $\left(\prod_{j=1}^{k}\sigma_{j}(y)\right)\;\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\,,$
where $A_{\alpha,i}(y):=D\varphi(\psi_{\alpha}^{i}(y))U(y)$, and $(\mathbb{I}_{n,k})_{l,m}:=\delta_{l,m}.$ Using the Cauchy-Binet formula [60, p. 68], one has
$\mathcal{H}^{k}(\varphi(S_{i}))=\sum_{\alpha}\int_{V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\rho_{\alpha}^{i}(\psi_{\alpha}^{i}(y))\left(\prod_{j=1}^{k}\sigma_{j}(y)\right)\;\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\mathrm{d}\mathcal{H}^{k}(y)$
$\leq$ $\sup_{\alpha}\left\|\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\right\|_{L^{\infty}(V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i}))}\cdot\sum_{\alpha}\int_{V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\rho_{\alpha}^{i}(\psi_{\alpha}^{i}(y))\left(\prod_{j=1}^{k}\sigma_{j}(y)\right)\mathrm{d}\mathcal{H}^{k}(y)$
$=$ $\sup_{\alpha}\left\|\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\right\|_{L^{\infty}(V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i}))}\cdot\sum_{\alpha}\int_{V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\rho_{\alpha}^{i}(\psi_{\alpha}^{i}(y))\llbracket D\psi_{\alpha}^{i}(y)\rrbracket\mathrm{d}\mathcal{H}^{k}(y)$
$=$ $\sup_{\alpha}\left\|\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\right\|_{L^{\infty}(V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i}))}\cdot\sum_{\alpha}\int_{U_{\alpha}^{i}\cap K_{i}}\rho_{\alpha}^{i}(x)\mathrm{d}\mathcal{H}^{k}(x)$
$=$ $\sup_{\alpha}\left\|\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\right\|_{L^{\infty}(V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i}))}\cdot\mathcal{H}^{k}(K_{i})$
(17) $=$ $\sup_{\alpha}\left\|\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\right\|_{L^{\infty}(V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i}))}\cdot\mathcal{H}^{k}(S_{i})\,.$
If $A$ is an $l\times m$ matrix with $l\geq m$ having singular values $\sigma_{1}\geq\dots\geq\sigma_{m}$, then the singular values $\sigma_{1}^{(1)}\geq\dots\geq\sigma_{m-1}^{(1)}$ of the matrix $A^{(1)}$ obtained removing a column from $A$ satisfy
$\sigma_{1}\geq\sigma_{1}^{(1)}\geq\sigma_{2}\geq\sigma_{2}^{(1)}\geq\cdots\geq\sigma_{m-1}\geq\sigma_{m-1}^{(1)}\geq\sigma_{m}\,,$
see, e.g., [39, Thm. 7.3.9]. Taking $l,m=n$, and iterating $n-k$ times the interlacing inequalities of (3.1.1) (removing the last column at each step), we can estimate each of the sigular values $\sigma_{j}^{(n-k)}$ of $A^{(n-k)}:=A\mathbb{I}_{n,k}$ by the corresponding singular value of $A$, i.e.,
$\sigma_{j}^{(n-k)}\leq\sigma_{j},\;\;j=1,2,\dots,k\,.$
Applying this result to $A_{\alpha,i}(y)$ for any $y\in V_{\alpha}^{i}$, the $j$-th singular value of $A_{\alpha,i}^{(n-k)}$ is bounded from above by the $j$-th singular value of $A_{\alpha,i}(y):=D\varphi(\psi_{\alpha}^{i}(y))U(y)$, which equals the $j$-th singular value $\sigma_{j}^{(\varphi)}(\psi_{\alpha}^{i}(y))$ of $D\varphi(\psi_{\alpha}^{i}(y))$ since $U(y)$ is orthogonal. Thus
(18) $\max_{y\in V_{\alpha}^{i}\cap\phi_{\alpha}^{i}(K_{i})}\sqrt{\det(\mathbb{I}_{n,k}^{\top}A_{\alpha,i}^{\top}(y)A_{\alpha,i}(y)\mathbb{I}_{n,k})}\leq\max_{x\in U_{\alpha}^{i}\cap K_{i}}\prod_{j=1}^{k}\sigma_{j}^{(\varphi)}(x),\quad\forall i,\;\forall\alpha\,.$
The combination of (3.1.1) with (18) proves
$\mathcal{H}^{k}(\varphi(S_{i}))\leq\left(\max_{x\in U_{\alpha}^{i}\cap K_{i}}\prod_{j=1}^{k}\sigma_{j}^{(\varphi)}(x)\right)\mathcal{H}^{k}(S_{i}),$
whence, summing over $i$, the rightmost inequality in (16) follows. The leftmost one can be obtained repeating the same reasoning, replacing $\varphi$ by $\varphi^{-1}$ and using the Inverse Function Theorem [38, Thm. 1.1.7]. ∎
Remark 7.
One may wonder if the assumption on the regularity of $\varphi$ in Lemma 2 can be relaxed to Lipschitz regularity on $\varphi$ and its inverse. The answer is negative, as in such a setting the differential of $\varphi$ may be not everywhere well defined on $S$.
Recall that $\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E}))$ denotes the Lebesgue constant associated with $\widehat{\mathcal{T}}$ (for the space spanned by $\mathscr{V}\widehat{E}$) and $\mathscr{L}({\mathcal{T}},\mathscr{V}(E))$ denotes the Lebesgue constant associated with ${\mathcal{T}}$ (for the space spanned by $\mathscr{V}(E)$). Lemma 1 and Lemma 2, together with the action of the pullback studied in Remark 3 and Remark 6, allow to relate $\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E}))$ and $\mathscr{L}({\mathcal{T}},\mathscr{V}(E))$.
Theorem 2.
Let $\varphi:\widehat{E}\to E$ be a $\mathscr{C}^{1}$ diffeomorphism. In the notation of Lemma 2, one has
(19) $\mathscr{L}({\mathcal{T}},\mathscr{V}(E))\leq\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})).$
Proof.
Plugging the bounds of Lemma 2 in Eq. (14) and exploiting Lemma 1, one finds
$\mathscr{L}({\mathcal{T}},\mathscr{V}(E))$ $=\sup_{\varphi(S)\in\mathcal{S}^{k}(E)}\sum_{i=1}^{N}|T_{\varphi(S)}(\omega_{i})|=\sup_{\varphi(S)\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(\varphi(S))}\sum_{i=1}^{N}\mathcal{H}^{k}(\varphi(S_{i}))\left|\int_{\varphi(S)}\omega_{i}\right|$
$\leq\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\frac{1}{\mathcal{H}^{k}(S)\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\sum_{i=1}^{N}\mathcal{H}^{k}(S_{i})\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}\left|\int_{S}\varphi^{*}\omega_{i}\right|$
$=\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\frac{1}{\mathcal{H}^{k}(S)}\sum_{i=1}^{N}\mathcal{H}^{k}(S_{i})\left|\int_{S}\varphi^{*}\omega_{i}\right|$
$=\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\sum_{i=1}^{N}|T_{S}(\varphi^{*}\omega_{i})|$
$=\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})),$
which shows the claim. ∎
When $\varphi$ is an affine mapping one gets a strong simplification of Lemma 2. Since the differential of $\varphi$ is the constant matrix $D\varphi=A$, in such a framework the estimate (16) reads:
(20) $\left(\prod_{j=n-k+1}^{n}\sigma_{j}\right)\mathcal{H}^{k}(S)\leq\mathcal{H}^{k}(\varphi(S))\leq\left(\prod_{j=1}^{k}\sigma_{j}\right)\mathcal{H}^{k}(S).$
Consequently, under this assumption Theorem 2 may be stated as follows.
Corollary 1.
Let $\varphi:\widehat{E}\to E$ be an affine mapping with linear part $A$. Let $\sigma_{1}\geq\ldots\geq\sigma_{n}$ be the singular values of $A$. Then
(21) $\mathscr{L}({\mathcal{T}},\mathscr{V}(E))\leq\frac{\prod_{j=1}^{k}\sigma_{j}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})).$
Observe that, when $\widehat{E}$ is a simplex, Corollary 1 reduces to Example 5. In the literature, similar results have been proved for $k=1$ (see [2, Lem. 2 and Prop. 1]) and for a general $k$ (see [18, Prop. 3.21]) under stricter assumptions. These results exploit different techniques, which for $k>1$ in turn yield the weaker result
(22) $\mathscr{L}({\mathcal{T}},\mathscr{V}(E))\leq\chi^{k}(A)\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})),$
the symbol $\chi$ denoting the conditioning $\chi:=\sigma_{1}/\sigma_{n}$ of the matrix $A$. It is immediate to deduce that Eq. (19) implies Eq. (22), since
$\mathscr{L}({\mathcal{T}},\mathscr{V}(E))\leq\frac{\prod_{j=1}^{k}\sigma_{j}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E}))\leq\frac{\sigma_{1}^{k}}{\sigma_{n}^{k}}\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E}))=\chi^{k}(A)\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})).$
Further, Eq. (19) shows that Eq. (22) is pessimistic for $k>\lfloor\frac{n}{2}\rfloor$. In fact, assume that $k>\lfloor\frac{n}{2}\rfloor$. Then $\sigma_{k}$ appears both at the numerator and the denominator of the right hand side of Eq. (19), and hence cancels out. In turn, we immediately deduce that Eq. (22) is improved by
$\mathscr{L}({\mathcal{T}},\mathscr{V}(E))\leq\min\left\{\chi^{k}(A),\chi^{n-k}(A)\right\}\mathscr{L}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})).$
Notice also that
$\min\left\{\chi^{k}(A),\chi^{n-k}(A)\right\}=\begin{cases}\chi^{k}(A)\quad&\text{if }k\leq\lfloor\frac{n}{2}\rfloor,\\ \chi^{n-k}(A)\quad&\text{if }k>\lfloor\frac{n}{2}\rfloor.\end{cases}$
Remark 8.
Theorem 2 may be reversed to show that the Lebesgue constant (14) does not depend on the supporting domain $E$ only for $k=0$ and $k=n$. While the nodal claim is known, see, e.g., [35], that for top dimensional forms has been explicitly shown only for $n=1$, see [20]. In all other cases, the dependence has been observed numerically, see [2, §6] and [18, §4.2.1]. Problem-oriented bounds may also be obtained, see [2, Rmk. 1].
3.2. Approximation: fitting a family of integral currents
In contrast with Section 3.1, suppose to be given a $\mathscr{V}(E)$-determining, but not necessarily unisolvent, set $\mathcal{T}:=\{T_{1},\ldots,T_{M}\}\subset\mathscr{A}^{k}(E)$. Given any set of positive weights $\{w_{1},\ldots,w_{M}\}$ we consider the discrete least squares projection $P$ defined in (10). We address the problem of computing $\|P\|_{\mathrm{op}}$.
In the context of discrete least squares, the role of the Lebesgue constant is played by the following quantity:
(23) $\mathscr{M}(\mathcal{T},\mathscr{V}(E)):=\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M}\left|\sum_{h=1}^{N}T_{i}(\eta_{h})T(\eta_{h})\right|w_{i},$
where $\{\eta_{h}\}_{h=1,\dots,N}$ is any orthonormal basis of $(\mathscr{V},(\cdot,\cdot)_{\mathcal{T},w})$. We remark that (23) does not depend on the particular choice of the orthonormal basis $\{\eta_{1},\dots,\eta_{N}\}.$ For, pick any other orthonormal basis $\{\theta_{1},\dots,\theta_{N}\}$, where, for some orthogonal matrix $C$, we have $\theta_{k}=\sum_{h=1}^{N}C_{h,k}\eta_{h}$, then directly compute the right hand side of (23) with $\eta_{h}$ replaced by $\theta_{h}$ and use the orthogonality of $C$.
The characterization for $P$ is analogous to that obtained for $\Pi$ in Proposition 1 and Theorem 1.
Theorem 3.
Let $\mathcal{T}=\{T_{1},\dots,T_{M}\}\subset\mathscr{A}^{k}(E)$ be a $\mathscr{V}$-determining subset and $\{w_{1},\dots,w_{M}\}$ be positive weights $w_{i}\geq 0$ for all $i$. Then the discrete least squares projection $P$ defined in (10) satisfies
(24) $\|P\|_{\mathrm{op}}\leq\mathscr{M}\,.$
Further, equality in Eq. (24) is attained in the case $\mathcal{H}^{k}(\operatorname*{supp}T_{i}\cap\operatorname*{supp}T_{j})=0$ for $i\neq j$.
Proof.
Since $(\mathscr{V},(\cdot,\cdot)_{\mathcal{T},w})$ is a $N$-dimensional Hilbert space continuously embedded in the Banach space $\mathscr{D}^{k}_{0}$, any zero order current $T$ admits a Riesz representer $K_{T}$. Indeed, picking any orthonormal basis $\{\eta_{1},\dots,\eta_{N}\}$ of $\mathscr{V}$, for any $\omega\in\mathscr{V}$ we can write
$T(\omega)=T\left(\sum_{h=1}^{N}(\omega,\eta_{h})_{\mathcal{T},w}\eta_{h}\right)=\sum_{h=1}^{N}(\omega,\eta_{h})_{\mathcal{T},w}T(\eta_{h})=\left(\omega,\sum_{h=1}^{N}T(\eta_{h})\eta_{h}\right)_{\mathcal{T},w}=:(\omega,K_{T})_{\mathcal{T},w}.$
By the orthogonality of the projection operator $P$ it follows that $(\omega,K_{T})_{\mathcal{T},w}=(P\omega,K_{T})_{\mathcal{T},w}$ for any $\omega\in\Lambda^{k}$. Thus we can write
$\|P\omega\|_{0}=$ $\sup_{T\in\mathscr{A}^{k}(E)}|T(P\omega)|=\sup_{T\in\mathscr{A}^{k}(E)}|(\omega,K_{T})_{\mathcal{T},w}|=\sup_{T\in\mathscr{A}^{k}(E)}\left|\sum_{i=1}^{M}T_{i}(\omega)T_{i}(K_{T})w_{i}\right|$
$\leq$ $\max_{j\in\{1,\dots,M\}}|T_{j}(\omega)|\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M}|T_{i}(K_{T})|w_{i}\leq\|\omega\|_{0}\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M}|T_{i}(K_{T})|w_{i}$
$=$ $\|\omega\|_{0}\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M}\left|\sum_{h=1}^{N}T_{i}(\eta_{h})T(\eta_{h})\right|w_{i}\,.$
Hence we have
$\|P\|_{\mathrm{op}}=\sup_{\|\omega\|_{0}=1}\|P\omega\|_{0}\leq\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M}\left|\sum_{h=1}^{N}T_{i}(\eta_{h})T(\eta_{h})\right|w_{i}.$
When $\mathcal{H}^{k}(\operatorname*{supp}T_{i}\cap\operatorname*{supp}T_{j})=0$ for $i\neq j$, one can follow the lines of the proof of Theorem 1 to show that, for any $T\in\mathscr{A}^{k}(E)$ and $\varepsilon>0$, there exists $\theta_{\varepsilon,T}\in\mathscr{D}_{0}^{k}(E)$ such that, for any $i=1,\dots,M$, has $\|\cdot\|_{0}$-norm equal to one ($\|\theta_{\varepsilon,T}\|_{0}=1$), is bounded away from zero on $\mathcal{T}$ ($|T_{i}(\theta_{\varepsilon,T})|\geq\left(1-\frac{\varepsilon}{\mathscr{M}}\right)\|\theta_{\varepsilon}\|_{0}$), and has a prescribed sign ($\operatorname*{sgn}T_{i}(\theta_{\varepsilon,T})=\operatorname*{sgn}T_{i}(K_{T})$). Hence $T_{i}(\theta_{\varepsilon,T})T_{i}(K_{T})=|T_{i}(\theta_{\varepsilon,T})||T_{i}(K_{T})|$ and we can compute
$\|P\|_{\mathrm{op}}$ $=\sup_{\omega\in\mathscr{D}_{0}^{k}(E)\,\|\omega\|_{0}=1}\sup_{T\in\mathscr{A}^{k}(E)}\left|\sum_{i=1}^{M}T_{i}(\omega)T_{i}(K_{T})w_{i}\right|\geq\left|\sum_{i=1}^{M}T_{i}(\theta_{\varepsilon,T})T_{i}(K_{T})w_{i}\right|$
$=\sum_{i=1}^{M}|T_{i}(\theta_{\varepsilon,T})||T_{i}(K_{T})|w_{i}\geq\left(1-\frac{\varepsilon}{\mathscr{M}}\right)\sum_{i=1}^{M}|T_{i}(K_{T})|w_{i}.$
Taking the supremum over $T\in\mathscr{A}^{k}(E)$ we obtain $\|P\|_{\mathrm{op}}\geq\mathscr{M}-\varepsilon$ and by arbitrariness of $\varepsilon>0$ we can conclude $\|P\|_{\mathrm{op}}=\mathscr{M}.$
Remark 9.
When $M=N$ and $w_{i}=1$ for all $i=1,\dots,N$, we recover the interpolation case. Indeed, since the Lagrange basis satisfies $T_{i}(\eta_{j})=\delta_{i,j}$, it is an orthonormal basis with respect to the scalar product induced by $\mathcal{T}$ and $w$. Hence Eq. (24) reads $\|P\|_{\mathrm{op}}\leq\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{N}|T(\eta_{i})|=\mathscr{L}.$
3.2.1. An $\ell^{2}$-type estimate of $\|P\|_{\mathrm{op}}$
In the proof of Theorem 3, we essentially used a $(1,\infty)$-Hölder inequality. It is also interesting to repeat the same computation exploiting a different choice of conjugate exponents. In particular, we can also estimate $\|P\|_{\mathrm{op}}$ by an $\ell^{2}$ technique. Performing the computations, one finds
$\|P\omega\|_{0}=$ $\sup_{T\in\mathscr{A}^{k}(E)}|T(P\omega)|=\sup_{T\in\mathscr{A}^{k}(E)}|\sum_{h=1}^{N}T(\eta_{h})(\eta_{h},\omega)_{\mathcal{T},w}|=\sup_{T\in\mathscr{A}^{k}(E)}\left|\left(\sum_{h=1}^{N}T(\eta_{h})\eta_{h},\omega\right)_{\mathcal{T},w}\right|$
$=$ $\sup_{T\in\mathscr{A}^{k}(E)}\left|(K_{T},\omega)_{\mathcal{T},w}\right|\leq\|K_{T}\|_{\mathcal{T},w}\|\omega\|_{\mathcal{T},w}\leq\max_{i}|T_{i}(\omega)|\left(\sum_{i=1}^{M}w_{i}\right)^{1/2}\|K_{T}\|_{\mathcal{T},w}$
$\leq$ $\|\omega\|_{0}\left(\sum_{i=1}^{M}w_{i}\right)^{1/2}\|K_{T}\|_{\mathcal{T},w}=\|\omega\|_{0}\left(\sum_{i=1}^{M}w_{i}\right)^{1/2}\left(\sum_{h=1}^{N}|T(\eta_{h})|^{2}\right)^{1/2}\,.$
Thus
$\|P\|_{\mathrm{op}}=\sup_{\omega\in\mathscr{D}_{0}^{k}(E)\setminus 0}\frac{\|P\omega\|_{0}}{\|\omega\|_{0}}\leq\sup_{T\in\mathscr{A}^{k}(E)}\left(\sum_{h=1}^{N}|T(\eta_{h})|^{2}\right)^{1/2}\left(\sum_{i=1}^{M}w_{i}\right)^{1/2}\,,$
and, in the normalized case $\sum_{i=1}^{M}w_{i}=1$, this reduces to
$\|P\|_{\mathrm{op}}\leq\sup_{T\in\mathscr{A}^{k}(E)}\left(\sum_{h=1}^{N}|T(\eta_{h})|^{2}\right)^{1/2}=:\mathscr{M}_{2}\,.$
It is worth noticing that $\mathscr{M}_{2}$ is indeed the best comparability constant between the $\|\cdot\|_{0}$ and the $\|\cdot\|_{\mathcal{T},w}$ norms on the space $\mathscr{V}$. In the case $\sum_{i=1}^{M}w_{i}=1$, we can further write
$\sup_{\omega\in\mathscr{V}}\left(\frac{\|\omega\|_{0}^{2}}{\|\omega\|_{\mathcal{T},w}^{2}}\right)^{1/2}\textbf{Fede: }{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}=}\sup_{\omega\in\mathscr{V}}\sup_{T\in\mathscr{A}^{k}(E)}\left(\frac{\sum_{h=1}^{N}|T(\eta_{h})|^{2}\;\sum_{h=1}^{N}(\eta_{h},\omega)_{\mathcal{T},w}^{2}}{\|\omega\|_{\mathcal{T},w}^{2}}\right)^{1/2}\leq\mathscr{M}_{2},$
where equality is attained for $\omega=(\sum_{h=1}^{N}T(\eta_{h})\eta_{h})/(\sum_{h=1}^{N}|T(\eta_{h})|^{2})^{1/2}.$
Remark 10.
In the case of nodal fitting (i.e. when functions and pointwise evaluations are involved) the quantity $\mathscr{M}_{2}$ corresponds to the diagonal of the reproducing kernel and often termed Bergman function. This framework is strictly connected with the theory of optimal experimental design, see, e.g., [41].
3.2.2. Dependence of $\mathscr{M}$ on the reference domain
In Section 3.1.1 we have seen that the the Lebesgue constant $\mathscr{L}$ depends on the reference domain $\widehat{E}$. When the reference domain is mapped via a $\mathscr{C}^{1}$ diffeomorphism, Theorem 2 bounds the change of $\mathscr{L}$. We now show that such a bound applies also to $\mathscr{M}$ under the same hypotheses. To this end, let us notice that for positive weights we have a scalar product on $\widehat{E}:=\varphi(E)$ as in Eq. (9) that reads
$(\widehat{\omega},\widehat{\eta})_{\widehat{\mathcal{T}},\widehat{w}}:=\sum_{i=1}^{M}\widehat{w}_{i}\widehat{T}_{i}(\omega)\widehat{T}_{i}(\theta)\,,$
hats stressing that objects are associated with $\widehat{E}$. Defining $\eta_{i}:=\varphi^{*}(\widehat{\eta}_{i})$ for $i=1,\ldots,N$, and the weights $\widehat{w}_{h}:=\left(\frac{\mathscr{H}^{k}\left(\varphi(S_{h})\right)}{\mathscr{H}^{k}(S_{h})}\right)^{2}w_{h}$ for $h=1,\ldots,M$, we may prove the following.
Lemma 3.
Let $\{T_{1},\ldots,T_{M}\}$ be a determining set for $\mathscr{V}(E)$. Then $\{\eta_{1},\ldots,\eta_{N}\}$ is an orthonormal basis for $\mathscr{V}(E)$ with respect to the weights $\{w_{1},\ldots,w_{M}\}$ if and only if $\{\widehat{\eta}_{1},\ldots,\widehat{\eta}_{N}\}$ is an orthonormal basis for $\mathscr{V}(\widehat{E})$ with respect to weights $\left\{\widehat{w}_{1},\ldots,\widehat{w}_{M}\right\}$.
Proof.
Unwinding the definition of the scalar product introduced in Eq. (9) and transforming averaging currents as in Remark 3, one finds
$\delta_{i,j}$ $=(\widehat{\eta}_{i},\widehat{\eta}_{j})_{\widehat{\mathcal{T}},\widehat{w}}=\sum_{h=1}^{M}\widehat{w}_{h}\widehat{T}_{h}(\widehat{\eta}_{i})T_{h}(\widehat{\eta}_{j})=\sum_{h=1}^{M}\widehat{w}_{h}\frac{1}{\mathcal{H}^{k}(\varphi(S_{h}))}\int_{\varphi(S_{h})}\widehat{\eta}_{i}\frac{1}{\mathcal{H}^{k}(\varphi(S_{h}))}\int_{\varphi(S_{h})}\widehat{\eta}_{j}$
$=\sum_{h=1}^{M}w_{h}\left(\frac{\mathscr{H}^{k}\left(\varphi(S_{h})\right)}{\mathscr{H}^{k}(S_{h})}\right)^{2}\frac{1}{\mathcal{H}^{k}(\varphi(S_{h}))}\int_{S_{h}}\varphi^{*}(\widehat{\eta}_{i})\frac{1}{\mathcal{H}^{k}(\varphi(S_{h}))}\int_{S_{h}}\varphi^{*}(\widehat{\eta}_{j})$
$=\sum_{h=1}^{M}w_{h}\frac{1}{\mathcal{H}^{k}(S_{h})}\int_{S_{h}}\eta_{i}\frac{1}{\mathcal{H}^{k}(S_{h})}\int_{S_{h}}\eta_{j}=(\eta_{i},\eta_{j})_{\mathcal{T},w},$
which proves the claim. ∎
The above result, together with Lemma 2, allows us to prove the counterpart of Theorem 2.
Proposition 2.
Let $\varphi:\widehat{E}\to E$ be a $\mathscr{C}^{1}$ diffeomorphism. In the notation of Lemma 2, one has
(25) $\mathscr{M}(\mathcal{T},\mathscr{V}(E))\leq\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\mathscr{M}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})).$
Proof.
Expanding Eq. (23) and using Lemma 3, one computes:
$\mathscr{M}(\mathcal{T},\mathscr{V}(E))$ $=\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M}\left|\sum_{h=1}^{N}w_{i}T_{i}(\eta_{h})T(\eta_{h})\right|$
$=\sup_{\varphi(S)\in\mathcal{S}^{k}(E)}\sum_{i=1}^{M}\left|\sum_{h=1}^{N}w_{i}\frac{1}{\mathcal{H}^{k}(\varphi(S_{i}))}\int_{\varphi(S_{i})}\eta_{h}\frac{1}{\mathcal{H}^{k}(\varphi(S))}\int_{\varphi(S)}\eta_{h}\right|$
$=\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\frac{1}{\mathcal{H}^{k}(\varphi(S))}\sum_{i=1}^{M}w_{i}\frac{1}{\mathcal{H}^{k}(\varphi(S_{i}))}\left|\sum_{h=1}^{N}\int_{S_{i}}\varphi^{*}\eta_{h}\int_{S}\varphi^{*}\eta_{h}\right|$
$=\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\frac{1}{\mathcal{H}^{k}(\varphi(S))}\sum_{i=1}^{M}\widehat{w}_{i}\frac{1}{\mathcal{H}^{k}(\varphi(S_{i}))}\left(\frac{\mathcal{H}^{k}(\varphi(S_{i}))}{\mathcal{H}^{k}(S_{i})}\right)^{2}\left|\sum_{h=1}^{N}\int_{S_{i}}\varphi^{*}\eta_{h}\int_{S}\varphi^{*}\eta_{h}\right|$
$=\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\frac{\mathcal{H}^{k}(S)}{\mathscr{(}H^{k}\varphi(S))}\sum_{i=1}^{M}\widehat{w}_{i}\frac{\mathcal{H}^{k}(\varphi(S_{i}))}{\mathcal{H}^{k}(S_{i})}\left|\sum_{h=1}^{N}\frac{1}{\mathcal{H}^{k}(S_{i})}\int_{S_{i}}\varphi^{*}\eta_{h}\frac{1}{\mathcal{H}^{k}(S)}\int_{S}\varphi^{*}\eta_{h}\right|$
$\leq\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\sup_{S\in\mathcal{S}^{k}(\widehat{E})}\sum_{i=1}^{M}\widehat{w}_{i}\left|\sum_{h=1}^{N}\frac{1}{\mathcal{H}^{k}S_{i}}\int_{S_{i}}\varphi^{*}\eta_{h}\frac{1}{\mathcal{H}^{k}S}\int_{S}\varphi^{*}\eta_{h}\right|$
$=\frac{\left\|\prod_{j=1}^{k}\sigma^{(\varphi)}_{j}\right\|_{\infty}}{\left(\left\|\prod_{j=n-k+1}^{n}(\sigma^{(\varphi)}_{j})^{-1}\right\|_{\infty}\right)^{-1}}\mathscr{M}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})),$
where the only inequality in the above chain follows from Lemma 2. The claim is proved. ∎
As in the case of interpolation, a strong simplification is obtained for $\varphi$ being an affine mapping. In such a case, Proposition 2 reduces to the following easy statement.
Corollary 2.
Let $\varphi:\widehat{E}\to E$ be an affine mapping with linear part $A$. Let $\sigma_{1}\geq\ldots\geq\sigma_{n}$ be the singular values of $A$. Then
(26) $\mathscr{M}(\mathcal{T},\mathscr{V}(E))\leq\frac{\prod_{j=1}^{k}\sigma_{j}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\mathscr{M}(\widehat{\mathcal{T}},\mathscr{V}(\widehat{E})).$
4. Admissible integral $k$-meshes
Given a Banach space $(\mathscr{F},\|\cdot\|)$ and a set of continuous linear functionals $\mathcal{T}:=\{T_{\alpha}\}$ on $\mathscr{F}$, we term $\mathcal{T}$ a norming set for a given subspace $\mathscr{V}$ of $\mathscr{F}$, if there exists $C<+\infty$ such that
(27) $\|f\|\leq C\max_{\alpha}|T_{\alpha}(f)|,\quad\forall f\in\mathscr{V}.$
Finite norming sets for polynomial spaces restricted to a given compact set are of leading interest in the design of approximation strategies such as interpolation, discrete least squares, and approximate optimal designs. Indeed, in [24] it was first observed that, for the case $\mathscr{F}=(\mathscr{C}^{0}(E),\|\cdot\|_{E})$ and $\mathscr{V}=\mathscr{P}_{r}$, the combination of property (27) with a bound on the cardinality of $\mathcal{T}$ leads to a number of remarkable consequences. This motivated the definition of admissible polynomial meshes as sequences $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ of finite subsets of $E$ such that
$\limsup_{r\to+\infty}\operatorname*{Card}(\mathcal{T}^{(r)})^{1/r}\leq 1$
(28) $C:=\sup_{r\in\mathbb{N}}\sup_{p\in\mathscr{P}_{r}\setminus\{0\}}\frac{\max_{x\in E}|p(x)|}{\max_{x\in\mathcal{T}^{(r)}}|p(x)|}<+\infty\,.$
The quantity $C$ in (28) is termed constant of the mesh. Such a definition can be weakened, defining
$C_{r}:=\sup_{p\in\mathscr{P}_{r}\setminus\{0\}}\frac{\max_{x\in E}|p(x)|}{\max_{x\in\mathcal{T}^{(r)}}|p(x)|}$
and allowing such $C_{r}$ to grow subexponentially with respect to $r$, i.e., $\limsup_{r\to+\infty}C_{r}^{1/r}\leq 1$. This generalization lead to the concept of weakly admissible polynomial meshes.
Admissible polynomial meshes found significant applications in the construction of quasi-optimal interpolation arrays and good discrete least squares projection [13, 12], the analysis of asymptotic behaviour of multivariate orthogonal polynomials and pluripotential theory [52], quadrature and optimal experimental designs [53, 16], and polynomial optimization [55]. For, the algorithmic construction of admissible polynomial meshes has been extensively studied and borrowed various techniques and tools from classical polynomial inequalities, convex geometry, and pluripotential theory [44, 45, 50, 54].
4.1. Integral meshes
The classical theory of admissible meshes has been designed around the uniform norm and considering pointwise evaluation functionals as sampling operators. These objects fit into the case $k=0$ of the present framework, as shown in Example 1. Drawing inspiration from the nodal case, the aim of the present section is to generalize the definition of admissible polynomial meshes to all cases $k\in\{1,\dots,n\}$. As one may expect, the most impactful changes concern the replacement of the uniform norm by $\|\cdot\|_{0}$, and the use of integral averages as sampling operators.
Definition 2 (Integral $k$-mesh).
Let $E\subset\mathbb{R}^{n}$ be a compact set with non-empty interior. For $k\in\{0,1,\dots,n\}$, we term integral $k$-mesh a sequence $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ of finite subsets $\mathcal{T}^{(r)}=\{T^{(s,r)}\}_{s=1,\dots,M(r)}$ of $\mathscr{A}^{k}(E)$, such that, for any $r\in\mathbb{N}$ and any $s=1,2,\dots,M(r)$, there exist $x^{(s,r)}\in E$, $A^{(s,r)}\in M_{n,k}(\mathbb{R}^{n})$ with $\operatorname*{Rank}(A^{(s,r)})=k$, a $\mathcal{L}^{k}$-measurable set $\Omega^{(s,r)}\subset\mathbb{R}^{k}$ with $0<\mathcal{L}^{k}(\Omega^{(s,r)})<+\infty$:
$T^{(s,r)}(\omega)=T_{S^{(s,r)}}(\omega)=\frac{1}{\mathcal{H}^{k}(S^{(s,r)})}\int_{S^{(s,r)}}\langle\omega;\sigma^{(s,r)}\rangle\mathrm{d}\mathcal{H}^{k}\,,\quad\forall\omega\in\mathscr{D}_{0}^{k}(E)\,,$
where $S^{(s,r)}$ is the $k$-rectifiable set $\{x^{(s,r)}+A^{(s,r)}y,y\in\Omega^{(s,r)}\}\subset E$ endowed by the orientation
$\sigma^{(s,r)}:=\frac{A_{:,1}^{(s,r)}\wedge\dots\wedge A_{:,k}^{(s,r)}}{|A_{:,1}^{(s,r)}\wedge\dots\wedge A_{:,k}^{(s,r)}|}\,,$
where we denoted by $A_{:,l}^{(s,r)}$ the $l$-th column of the matrix $A^{(s,r)}.$
By a slight abuse of notation and terminology, we can identify $\mathcal{T}^{(r)}$ with the set of oriented supports $\{(x^{1,r},A^{(1,r)},\Omega^{(1,r)}),\cdots,(x^{M(r),r},A^{(M(r),r)},\Omega^{(M(r),r)})\}$, and term the latter integral $k$-mesh as well.
Remark 11.
We decided to include in the Definition 2 the requirement of $S^{(s,r)}$ being a piece of an affine variety. This is mainly motivated from the perspective of applications and simplicity of coding. In particular, in this setting one has
$T_{S^{(s,r)}}(\omega)=$ $\frac{1}{\mathcal{H}^{k}(S^{(s,r)})}\int_{S^{(s,r)}}\langle\omega;\sigma^{(s,r)}\rangle\mathrm{d}\mathcal{H}^{k}$
$=$ $\frac{1}{\int_{\Omega^{(s,r)}}\llbracket A^{(s,r)}\rrbracket\mathrm{d}y}\int_{\Omega^{(s,r)}}\langle\omega(x^{(s,r)}+A^{(s,r)}y);\sigma^{(s,r)}\rangle\llbracket A^{(s,r)}\rrbracket\mathrm{d}y$
$=$ $\frac{1}{\mathcal{L}^{k}(\Omega^{(s,r)})}\int_{\Omega^{(s,r)}}\langle\omega(x^{(s,r)}+A^{(s,r)}y);\sigma^{(s,r)}\rangle\mathrm{d}y\,.$
The sequence $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ naturally induces a sequence of seminorms on $\mathscr{D}^{k}_{0}(E)$ defined by
$\|\omega\|_{\mathcal{T}^{(r)}}:=\max_{s\in\{1,\dots,M(r)\}}|T^{(s,r)}(\omega)|\;,\;\;\forall\omega\in\mathscr{D}_{0}^{k}(E)\,,$
which is used to extend the definition of admissible polynomial meshes to differential forms.
Definition 3 ((Weakly-)admissible integral $k$-mesh ).
Let $\{\mathscr{V}^{(r)}\}_{r\in\mathbb{N}}$ be an increasing sequence of linear subspaces of $\mathscr{D}_{0}^{k}(E).$ Let $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ be an integral $k$-mesh for the compact set $E\subset\mathbb{R}^{n}$. Then $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ is termed a $\{\mathscr{V}^{(r)}\}_{r\in\mathbb{N}}$-admissible integral $k$-mesh if, in the above notation,
$\limsup_{r\to+\infty}\left[\operatorname*{Card}(\mathcal{T}^{(r)})\right]^{1/r}\leq 1$
(29) $C:=\sup_{r\in\mathbb{N}}\;\sup\left\{\frac{\|\omega\|_{0}}{\|\omega\|_{\mathcal{T}^{(r)}}},\ \omega\in\mathscr{V}^{(r)}\setminus\{0\}\right\}<+\infty\,.$
If the property (29) is replaced by the weaker assumption
$\limsup_{r}C_{r}^{1/r}:=\limsup_{r}\left(\sup\left\{\frac{\|\omega\|_{0}}{\|\omega\|_{\mathcal{T}^{(r)}}},\ \omega\in\mathscr{V}^{(r)}\setminus\{0\}\right\}\right)^{1/r}\leq 1\,,$
then $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ is termed a $\{\mathscr{V}^{(r)}\}_{r\in\mathbb{N}}$-weakly admissible integral $k$-mesh.
In the following Lemma 4 we collect some basic properties of integral $k$-meshes, which we invite the reader to compare with [11, §4]. We point out that some of the features of weakly admissible (00-)meshes do not have natural counterparts in the context of integral $k$-meshes with $k>0$.
Lemma 4.
The following properties of integral $k$-meshes hold true:
Subspace, if $\{\mathcal{T}^{(r)}\}$ is a $\{\mathscr{V}^{r}\}$-admissible integral $k$-mesh and $\mathscr{W}^{r}\subset\mathscr{V}^{r}$ for any $r$, then $\{\mathcal{T}^{(r)}\}$ is a $\{\mathscr{W}^{r}\}$-admissible integral $k$-mesh as well;
Finite unions, let $E=\cup_{i=1}^{m}E_{i}$ and denote by $\mathscr{V}^{(r)}_{i}$ the space of the restrictions of the elements of $\mathscr{V}^{(r)}$ to $E_{i}$. If, for any $i=\{1,\dots.m\}$, $\{\mathcal{T}_{i}^{(r)}\}_{r\in\mathbb{N}}$ is a $\{\mathscr{V}^{(r)}_{i}\}_{r\in\mathbb{N}}$-admissible integral $k$-mesh for $E_{i}$, then $\{\cup_{i=1}^{m}\mathcal{T}^{(r)}_{i}\}_{r\in\mathbb{N}}$ is a $\{\mathscr{V}^{(r)}\}_{r\in\mathbb{N}}$-admissible integral $k$-mesh for $E$;
Good unisolvent triangular arrays in $\mathscr{A}^{k}(E)$, assume that $\operatorname*{Card}\mathcal{T}^{(r)}=\operatorname*{dim}\mathscr{V}^{(r)}$ for any $r$, and that $\limsup_{r}[\mathscr{L}(\mathcal{T}^{(r)},\mathscr{V}^{r})]^{1/r}\leq 1$. Then $\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ is termed a $\{\mathscr{V}^{(r)}\}_{r\in\mathbb{N}}$-weakly integral $k$-mesh;
Affine mappings, let $\widehat{\mathcal{T}}^{(r)}=\{T_{\widehat{S}_{1}},\dots,T_{\widehat{S}_{M(r)}}\}$ be an admissible integral $k$-mesh of constant $\widehat{C}$ for the compact set $\widehat{E}$. Let $\varphi:\widehat{E}\rightarrow E:=\varphi(\widehat{E})$ be a non degenerate affine mapping with linear part having singular values $\sigma_{1}\geq\ldots\geq\sigma_{n}.$ Then $\mathcal{T}^{(r)}:=\{T_{\varphi(\widehat{S}_{1})},\dots,T_{\varphi(\widehat{S}_{M(r)})}\}$ is an admissible integral $k$-mesh whose constant $C$ satisfies
$C\leq\frac{\prod_{j=1}^{k}\sigma_{j}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\widehat{C}\,.$
Proof.
The first three properties follows immediately from Definition 3. The fourth follows from Lemma 2 and the sampling inequality (29):
$\|\omega\|_{0}=$ $\sup_{S\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(S)}\left|\int_{S}\omega\right|=\sup_{\varphi(\widehat{S})\in\mathcal{S}^{k}(E)}\frac{1}{\mathcal{H}^{k}(\varphi(\widehat{S}))}\left|\int_{\varphi(\widehat{S})}\omega\right|=\sup_{\widehat{S}\in\mathcal{S}^{k}(\widehat{E})}\frac{1}{\mathcal{H}^{k}(\varphi(\widehat{S}))}\left|\int_{\widehat{S}}\varphi^{*}\omega\right|$
$=$ $\sup_{\widehat{S}\in\mathcal{S}^{k}(\widehat{E})}\frac{\mathcal{H}^{k}(\widehat{S})}{\mathcal{H}^{k}(\varphi(\widehat{S}))}|T_{\widehat{S}}(\varphi^{*}\omega)|\leq\sup_{\widehat{S}\in\mathcal{S}^{k}(\widehat{E})}\frac{1}{\prod_{j=n-k+1}^{n}\sigma_{j}}|T_{\widehat{S}}(\varphi^{*}\omega)|=\frac{1}{\prod_{j=n-k+1}^{n}\sigma_{j}}\|\varphi^{*}\omega\|_{0}$
$\leq$ $\frac{\widehat{C}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\|\varphi^{*}\omega\|_{\widehat{\mathcal{T}}^{(r)}}=\frac{\widehat{C}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\max_{i}|T_{\widehat{S}_{i}}(\varphi^{*}\omega)|$
$=$ $\frac{\widehat{C}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\max_{i}\frac{\mathcal{H}^{k}(\varphi(\widehat{S}_{i}))}{\mathcal{H}^{k}(\widehat{S}_{i})}|T_{\varphi(\widehat{S}_{i})}(\omega)|\leq\frac{\prod_{j=1}^{k}\sigma_{j}}{\prod_{j=n-k+1}^{n}\sigma_{j}}\widehat{C}\|\omega\|_{\mathcal{T}^{(r)}}\,.$
The claim is proved. ∎
The concept of admissibility for integral $k$-meshes given in Definition 3 applies to any increasing sequence $\{\mathscr{V}^{(r)}\}_{r\in\mathbb{N}}$ of subspaces of $\mathscr{D}_{0}^{k}(E)$, but clearly the construction of specific admissible meshes depends on the considered sequence. The next two subsections are devoted to the construction of admissible integral $k$-meshes for the polynomial case $\mathscr{V}^{r}(E)=\mathscr{P}_{r}\Lambda^{k}(E)$. We remark that, since $\mathscr{P}_{r}^{-}\Lambda^{k}(E)\subset\mathscr{P}_{r}\Lambda^{k}(E)$, any $\{\mathscr{P}_{r}\Lambda^{k}(E)\}$-admissible integral $k$-mesh is also a $\mathscr{P}_{r}^{-}\Lambda^{k}(E)$-admissible integral $k$-mesh due to the first property in Lemma 4.
4.2. Constructing $\mathscr{P}_{r}\Lambda^{k}$-admissible integral $k$-meshes by Markov inequality
The Markov inequality, here recalled in the forthcoming Eq. (30), is a classical tool in approximation theory. Because of its ductility, it has been investigated for long time and now presents broad scope connections with several branches of mathematics; for an account, we address the reader to [49, 5].
Let us recall that a polynomial determining compact set $E\subset\mathbb{R}^{n}$ admits a Markov inequality if there exist $C_{M}<+\infty$ (the Markov constant) and $\beta<+\infty$ (the Markov exponent) such that, for any polynomial $p$ of degree at most $r$, one has
(30) $\max_{x\in E}|\nabla p(x)|\leq C_{M}r^{\beta}\max_{x\in E}|p(x)|\,.$
As proved in [64], any fat convex body (i.e. $E=\overline{\operatorname*{int}E}$) admits a Markov inequality with exponent $\beta=2$ and constant $C_{M}$ equal to the reciprocal of the minimum distance of two supporting hyperplanes.
Theorem 4 (Fundamental estimate on convex bodies).
Let $E\subset\mathbb{R}^{n}$ be a convex fat set with Markov constant $C_{M}$. Let $\mathcal{T}=\{T^{(s)}\}_{s=1,\dots,M}\subset\mathscr{A}^{k}(E)$ be such that, for any $s=1,2,\dots,M$, there exist $x^{(s)}\in E$, $A^{(s)}\in M_{n,k}(\mathbb{R}^{n})$ with $\operatorname*{Rank}(A^{(s)})=k$, a $\mathcal{L}^{k}$-measurable set $\Omega^{(s)}\subset\mathbb{R}^{k}$ with $0<\mathcal{H}^{k}(\Omega^{(s)})<+\infty$, for which
$T^{(s)}(\omega)=T_{S^{(s)}}(\omega)=\frac{1}{\mathcal{H}^{k}(S^{(s)})}\int_{S^{(s)}}\langle\omega;\sigma^{(s)}\rangle\mathrm{d}\mathcal{H}^{k}\;,\;\;\forall\omega\in\mathscr{D}_{0}^{k}(E)\,,$
where $S^{(s)}$ is the $k$-rectifiable set $\{x^{(s)}+A^{(s)}y,y\in\Omega^{(s)}\}\subset E$ endowed by the orientation
$\sigma^{(s)}:=\frac{A_{:,1}^{(s)}\wedge\dots\wedge A_{:,k}^{(s)}}{|A_{:,1}^{(s)}\wedge\dots\wedge A_{:,k}^{(s,r)}|}\,.$
Let $r\in\mathbb{N}$ and assume that there exists $c_{1}<1$ such that, for any $x\in E$, we can find $s_{1},s_{2},\dots,s_{m}\in\{1,\dots,M\}$ such that
(31) $\sup_{\tau\in\Lambda_{k}}\Bigg{\{}\min_{a\in\mathbb{R}^{m}}\Big{\{}|a|_{1}:\sum_{j=1}^{m}a_{j}\sigma^{(s_{j})}=\tau\Big{\}},|\tau|=1,\tau\text{ simple}\Bigg{\}}=:c_{2}<+\infty$
(32) $\max_{j=1,\dots,m}\max_{z\in S^{s_{j}}}|x-z|\leq c_{1}\left(c_{2}C_{M}r^{2}\sqrt{\binom{n}{k}}\right)^{-1}\,.$
Then, for any $\omega\in\mathscr{P}_{r}\Lambda^{k}$ we have
(33) $\|\omega\|_{0}\leq\frac{c_{2}}{1-c_{1}}\;\max_{s=1,\dots,M}\left|T_{S^{s}}(\omega)\right|=:\frac{c_{2}}{1-c_{1}}\|\omega\|_{\mathcal{T}}.$
Proof.
Let $\omega\in\mathscr{P}_{r}\Lambda^{k}$ be any polynomial form. We prove that (33) holds for $\omega$ under the assumption that there exist a point $\bar{x}\in E$ and a simple $k$-vector $\bar{\tau}\in\Lambda_{k}$ such that
(34) $\max_{x\in E}|\omega(x)|^{*}=\langle\omega(\bar{x});\bar{\tau}\rangle,$
the general case will readily follow by approximation.
Let $m\in\mathbb{N}$, $s_{1},s_{2},\dots,s_{m}\in\{1,\dots,M\}$ and $c_{1}<1$ be such that
(35) $\max_{j=1,\dots,m}\max_{z\in S^{s_{j}}}|\bar{x}-z|\leq c_{1}\left(c_{2}C_{M}r^{2}\sqrt{\binom{n}{k}}\right)^{-1},$
and let $\bar{a}\in\mathbb{R}^{m}$ be any vector realizing
(36) $\min_{a\in\mathbb{R}^{m}}\Big{\{}|a|_{1}:\sum_{j=1}^{m}a_{j}\sigma^{(s_{j})}=\bar{\tau}\Big{\}}\,.$
Plugging (36) into (34), we compute
$\max_{x\in E}|\omega(x)|^{*}=\langle\omega(\bar{x});\bar{\tau}\rangle=\sum_{j=1}^{m}\bar{a}_{j}\langle\omega(\bar{x});\sigma^{(s_{j})}\rangle=\sum_{j=1}^{m}\bar{a}_{j}\frac{1}{\mathcal{H}^{k}(S^{s_{j}})}\int_{S^{s_{j}}}\langle\omega(\bar{x});\sigma^{(s_{j})}\rangle\mathrm{d}\mathcal{H}^{k}(x)$
$=$ $\sum_{j=1}^{m}\bar{a}_{j}\frac{1}{\mathcal{H}^{k}(S^{s_{j}})}\int_{S^{s_{j}}}\langle\omega(x);\sigma^{(s_{j})}\rangle\mathrm{d}\mathcal{H}^{k}(x)+\sum_{j=1}^{m}\bar{a}_{j}\frac{1}{\mathcal{H}^{k}(S^{s_{j}})}\int_{S^{s_{j}}}\langle\omega(\bar{x})-\omega(x);\sigma^{(s_{j})}\rangle\mathrm{d}\mathcal{H}^{k}(x)$
$\leq$ $|\bar{a}|_{1}\|\omega\|_{\mathcal{T}}+|\bar{a}|_{1}\max_{j=1,\dots,m}\frac{1}{\mathcal{H}^{k}(S^{s_{j}})}\int_{S^{s_{j}}}|\langle\omega(\bar{x})-\omega(x);\sigma^{(s_{j})}\rangle|\mathrm{d}\mathcal{H}^{k}(x)$
$\leq$ $c_{2}\left(\|\omega\|_{\mathcal{T}}+\max_{j=1,\dots,m}\frac{1}{\mathcal{H}^{k}(S^{s_{j}})}\int_{S^{s_{j}}}|\omega(\bar{x})-\omega(x)|\mathrm{d}\mathcal{H}^{k}(x)\right)$
$=$ $c_{2}\left(\|\omega\|_{\mathcal{T}}+\max_{j=1,\dots,m}\frac{1}{\llbracket A^{s_{j}}\rrbracket\mathcal{L}^{k}(\Omega^{s_{j}})}\int_{\Omega^{s_{j}}}|\omega(\bar{x})-\omega(x^{s_{j}}+A^{s_{j}}y)|\llbracket A^{s_{j}}\rrbracket\mathrm{d}y\right)$
$=$ $c_{2}\left(\|\omega\|_{\mathcal{T}}+\max_{j=1,\dots,m}\frac{1}{\llbracket A^{s_{j}}\rrbracket\mathcal{L}^{k}(\Omega^{s_{j}})}\int_{\Omega^{s_{j}}}\left({\sum_{\alpha}}^{\prime}(\nabla\omega_{\alpha}(\xi_{\alpha,j,y});\bar{x}-x^{s_{j}}-A^{s_{j}}y)^{2}\right)^{1/2}\llbracket A^{s_{j}}\rrbracket\mathrm{d}y\right)$
(37) $=$ $c_{2}\left(\|\omega\|_{\mathcal{T}}+\max_{j=1,\dots,m}\frac{1}{\mathcal{L}^{k}(\Omega^{s_{j}})}\int_{\Omega^{s_{j}}}\left({\sum_{\alpha}}^{\prime}(\nabla\omega_{\alpha}(\xi_{\alpha,j,y});\bar{x}-x^{s_{j}}-A^{s_{j}}y)^{2}\right)^{1/2}\mathrm{d}y\right)\,,$
where the point $\xi_{\alpha,j,y}$ lies in the segment $[\bar{x},x^{s_{j}}+A^{s_{j}}y]$. Notice that $\xi_{\alpha,j,y}\in E$ since $E$ is convex.
Using the Cauchy-Schwartz Inequality, the bound of Eq. (35), and the Markov Inequality (30), we obtain that, for each $y\in\Omega^{s_{j}}$ and $j\in\{1,\dots,m\}$, the following inequality holds:
$\left({\sum_{\alpha}}^{\prime}(\nabla\omega_{\alpha}(\xi_{\alpha,j,y}),\bar{x}-x^{s_{j}}-A^{s_{j}}y)^{2}\right)^{1/2}\leq\left({\sum_{\alpha}}^{\prime}|\nabla\omega_{\alpha}(\xi_{\alpha,j,y})|^{2}\right)^{1/2}|\bar{x}-x^{s_{j}}-A^{s_{j}}y|$
$\leq$ $\left({\sum_{\alpha}}^{\prime}|\nabla\omega_{\alpha}(\xi_{\alpha,j,y})|^{2}\right)^{1/2}\max_{j=1,\dots,m}\max_{z\in S^{s_{j}}}|\bar{x}-z|\leq\left({\sum_{\alpha}}^{\prime}|\nabla\omega_{\alpha}(\xi_{\alpha,j,y})|^{2}\right)^{1/2}c_{1}\left(c_{2}C_{M}r^{2}\sqrt{\binom{n}{k}}\right)^{-1}$
$\leq$ $\left({\sum_{\alpha}}^{\prime}C_{M}^{2}r^{4}\max_{x\in E}|\omega_{\alpha}(x)|^{2}\right)^{1/2}c_{1}\left(c_{2}C_{M}r^{2}\sqrt{\binom{n}{k}}\right)^{-1}=\left({\sum_{\alpha}}^{\prime}\max_{x\in E}|\omega_{\alpha}(x)|^{2}\right)^{1/2}c_{1}\left(c_{2}\sqrt{\binom{n}{k}}\right)^{-1}$
(38) $\leq$ $\sqrt{\binom{n}{k}}\max_{\alpha}\max_{x\in E}|\omega_{\alpha}(x)|c_{1}\left(c_{2}\sqrt{\binom{n}{k}}\right)^{-1}=\frac{c_{1}}{c_{2}}\max_{\alpha}\max_{x\in E}|\omega_{\alpha}(x)|\leq\frac{c_{1}}{c_{2}}\max_{x\in E}|\omega(x)|^{*}\,.$
Finally, using Eq. (38) to estimate the last term of Eq. (37), we obtain
$\max_{x\in E}|\omega(x)|^{*}\leq c_{2}\left(\|\omega\|_{\mathcal{T}}+\frac{c_{1}}{c_{2}}\max_{x\in E}|\omega(x)|^{*}\right).$
Since we are assuming $c_{1}<1$, Eq. (33) follows. ∎
Remark 12.
The convexity assumption of Theorem 4 strongly simplifies computations, but is not strictly necessary. The extension to more general compact sets satisfying a Markov inequality can be obtained by a suitable modification of the argument in the proof, following the lines of [24, Thm. 5].
4.2.1. An explicit example of integral $k$-mesh by Markov inequality
We show how Theorem 4 can be used to construct an integral $k$-mesh on a real fat convex body $E$.
Let $10<c_{1}<$ 1 and $r\in\mathbb{N}$. Consider a bounding box $Q:=[a_{1},b_{1}]\times[a_{2},b_{2}]\times\dots\times[a_{n},b_{n}]\supset E$ and a tessellation $\{Q_{i}\}_{i=1,\dots,\tilde{N}}$ of $Q$ made of coordinate $n$-dimensional cubes with side lengths $d$ bounded from above by
(39) $d:=\frac{c_{1}w(E)}{4\sqrt{n}\binom{n}{k}r^{2}},\;w(E):=\min_{v\in\mathbb{R}^{n}}\min\{\ell:|(x;v)|\leq\ell\;\forall x\in E\}.$
For any $i=1,\dots,m$ such that $K_{i}:=Q_{i}\cap\operatorname*{int}E\neq\emptyset$, pick $\tilde{x}_{i}\in E_{i}$ and define
$\{x^{(s,r)}\}_{s=1}^{M(r)}:=(\,\overbrace{\tilde{x}_{1},\ldots,\tilde{x}_{1}}^{\binom{n}{k}\text{ times}},\overbrace{\tilde{x}_{2},\ldots,\tilde{x}_{2}}^{\binom{n}{k}\text{ times}},\ldots,\overbrace{\tilde{x}_{m},\ldots,\tilde{x}_{m}}^{\binom{n}{k}\text{ times}}\,)\,.$
Pick
$0<\epsilon^{(s,r)}\leq\min\left(\operatorname*{dist}(x^{(s,r)},\partial E),\operatorname*{dist}(x^{(s,r)},\partial Q_{i(s)})\right),$
where $i(s)$ is the unique index such that $x^{(s,r)}\in Q_{i(s)}$, and set $\Omega^{(s,r)}\equiv\epsilon^{(s,r)}\Sigma_{k},\,\forall s=1,\dots,M(r).$ Given the enumeration $\alpha_{1},\alpha_{2},\dots,\alpha_{\binom{n}{k}}$ of the set of multi-indices $\alpha$ of length $k$ we set
$\{A^{1,r},A^{2,r},\dots,A^{\binom{n}{k},r},A^{\binom{n}{k}+1,r},\dots,A^{M(r),r}\}:=\{P^{\alpha_{1}},P^{\alpha_{2}},\dots,P^{\alpha_{\binom{n}{k}}},P^{\alpha_{1}},\dots,P^{\alpha_{\binom{n}{k}}}\},$
where $P^{\alpha}$ is the $n$ by $k$ matrix whose columns are the vectors $e_{\alpha(1)},\dots,e_{\alpha(k)}$ of the canonical basis of $\mathbb{R}^{n}$. Finally consider the collection
$\mathcal{T}^{(r)}:=\{(x^{(s,r)},\Omega^{(s,r)},A^{(s,r)}),s=1,\dots,M(r)\}.$
An example of this construction is reported in Figure 1. Note in particular that, since any $x\in E$ belongs to the closure of some $K_{i}$, we can find $\binom{n}{k}$ elements $(x^{(s_{j},r)},\Omega^{(s_{j},r)},A^{(s_{j},r)})$ of $\mathcal{T}^{(r)}$ such that
  1. a)
    $x^{(s_{j},r)}\equiv\tilde{x}_{i}$ for $j=1,2,\dots,\binom{n}{k}$;
  2. b)
    The set of $k$-vectors $\tau_{j}=\frac{A^{(s_{j},r)}_{:,1}\wedge A^{(s_{j},r)}_{:,2}\dots\wedge A^{(s_{j},r)}_{:,k}}{|A^{(s_{j},r)}_{:,1}\wedge A^{(s_{j},r)}_{:,2}\dots\wedge A^{(s_{j},r)}_{:,k}|}$ is an orthonormal basis of $\Lambda_{k}$;
  3. c)
    Setting $F^{(s_{j})}:=\{x^{(s_{j},r)}+A^{(s_{j},r)}y,y\in\Omega^{(s_{j},r)}\}$, we have $\max_{j=1,\dots,\binom{n}{k}}\max_{z\in\mathcal{F}^{(s_{j})}}|x-z|\leq\sqrt{n}d.$
Also, note that if we consider the canonical basis $\{e_{\alpha(1)}\wedge\dots\wedge e_{\alpha(k)}:\,|\alpha|=k\}$ of $\Lambda_{k}(\mathbb{R}^{n})$, then any unit simple $k$-vector $\tau$ can be written in the form $\tau=\sum^{\prime}_{|\alpha|=k}a_{\alpha}e_{\alpha(1)}\wedge\dots\wedge e_{\alpha(k)}$ with
$|a|_{1}\leq\sqrt{\binom{n}{k}}.$
Thus, due to b) above, we can assume (31) with $c_{2}=\sqrt{\binom{n}{k}}.$
On the other hand, we recall that the Markov constant of $E$ is bounded by $4/w(E)$ due to the classical result in [64]. Thus, combining (39) with property c) above, we have
$\max_{j=1,\dots,\binom{n}{k}}\max_{z\in\mathcal{F}^{(s_{j})}}|x-z|\leq\sqrt{n}d=\frac{c_{1}w(E)}{4\binom{n}{k}r^{2}}=\frac{c_{1}}{C_{M}r^{2}\binom{n}{k}},$
i.e., (32) holds as well.
Finally, we can give an asymptotic upper bound to $\operatorname*{Card}\mathcal{T}^{(r)}$ noticing that the number of $k$-faces of the tessellation that lie $E$ is smaller that the number of $k$-faces of the tessellation. The latter is at most $\binom{n}{k}$ times the number of cubes, thus
$\operatorname*{Card}\mathcal{T}^{(r)}\leq\binom{n}{k}\left(\frac{\mathrm{diam}(E)}{d}\right)^{n}=\binom{n}{k}^{n+1}\left(\frac{\mathrm{diam}(E)}{w(E)}\frac{4\sqrt{n}}{c_{1}}\right)^{n}r^{2n}=\mathcal{O}(r^{2n})\,.$
Since the hypotheses of Theorem 4 are fulfilled, we can claim the following result.
Proposition 3 (Admissible integral $k$-mesh on fat convex body).
The sequence $\mathcal{T}=\{\mathcal{T}^{(r)}\}_{r\in\mathbb{N}}$ defined in (4.2.1) is an admissible integral $k$-mesh for $E$ with constant $C:=\frac{\sqrt{\binom{n}{k}}}{1-c_{1}},$ i.e.
$\|\omega\|_{0}\leq\frac{\sqrt{\binom{n}{k}}}{1-c_{1}}\|\omega\|_{\mathcal{T}^{(r)}},\quad\forall\omega\in\mathscr{P}_{r}\Lambda^{k}(E).$
Figure 1. A visual representation of the construction of Proposition 3 in the case $n=2$, $k=1$. The thick line represents the boundary of the convex fat set $E\subset\mathbb{R}^{2}$. Gray squares are the squares $Q_{i}$’s of the considered coordinate tessellation that intersect $E$. In this example the points $\tilde{x}_{i}$ and the two columns of the matrices $A^{(s,r)}$ are randomly chosen. Note that in such a way there is no control on the size of the support of the constructed functionals.
Example 6 (An admissible integral $2$-mesh for the unit cube).
We can construct an admissible integral $2$-mesh for the unit cube $E:=[0,1]^{3}$ by slightly sharpening the construction of Proposition 3.
We consider a tessellation of $E$ made of coordinate cubes with side
$d=\frac{c_{1}w(E)}{2\sqrt{n}\binom{n}{k}r^{2}}=\frac{c_{1}}{2\sqrt{n}\binom{n}{k}r^{2}}\,,$
where $c_{1}<1$ is such that $d^{-1}\in\mathbb{N}$. For any $x\in E$ we can find a closed cube $Q$ of the tessellation for which $x\in Q$. Let $v$ be the closest vertex of $Q$ to $x$ and let $\mathcal{F}^{(1)},\mathcal{F}^{(2)},\mathcal{F}^{(3)}$ be the faces of $Q$ containing $v$. It is clear that the $2$-vectors $\tau^{(1)},\tau^{(2)},\tau^{(3)}$ representing any orientation of $\mathcal{F}^{(1)},\mathcal{F}^{(2)},\mathcal{F}^{(3)}$ form an orthonormal basis of $\Lambda_{k}$ (i.e., property b) above still holds). It is also easy to see that
$\max_{j=1,\dots,\binom{n}{k}}\max_{z\in\mathcal{F}^{(s_{j})}}|x-z|\leq\sqrt{n}d/2.$
Denote by $\mathcal{T}^{(r)}$ the set of integral averaging currents supported on all faces of the tessellation. By Theorem 4, we have
$\|\omega\|_{0}\leq\frac{\sqrt{\binom{n}{k}}}{1-c_{1}}\|\omega\|_{\mathcal{T}^{r}},\;\forall\omega\in\mathscr{P}_{r}\Lambda^{k}(E),$
that is, $\mathcal{T}^{(r)}$ is an integral $2$-mesh of constant $\sqrt{\binom{n}{k}}/(1-c_{1})$. The cardinality of this mesh is given by the number of faces in the tessellation, which equals to
$\operatorname*{Card}\mathcal{T}^{(r)}=\binom{3}{2}\frac{1}{d^{k}}\left(1-\frac{1}{d}\right)^{n-k}=\frac{324r^{4}(6\sqrt{3}r^{2}+c_{1})}{c_{1}^{3}}\sim\frac{1944}{c_{1}^{3}}r^{6}.$
Figure 2. Cardinalities of $\mathscr{P}_{r}\Lambda^{k}$-admissible integral $k$-meshes for the square (i.e., $n=2$) and the cube (i.e., $n=3$) constructed following Example 6 with $r$ varying from $1$ to $20$, and $k=1,\dots,n$, and $c_{1}=1/2$.
Remark 13.
The size of meshes constructed by the technology based on Theorem 4 grows very quickly, see Figure 2. Although this strategy is very flexible and can be easily adapted to many scenarios, meshes of such a large cardinality are not of practical use for most applications. This motivates the development of an alternative technique, which is carried in the next section.
4.3. Constructing $\mathscr{P}_{r}\Lambda^{k}$-admissible integral $k$-mesh by Baran inequality
The classical Bernstein Inequality on the interval $E:=[-1,1]$ states that, for any polynomial $p$ such that $|p(x)|\leq 1$ for any $x\in E$, one has
(40) $\frac{|p^{\prime}(x)|}{\sqrt{\max_{y\in E}|p(y)|^{2}-|p(x)|^{2}}}\leq\deg(p)\frac{1}{\sqrt{1-x^{2}}},\quad\forall x\in\operatorname*{int}E.$
This inequalilty has immediate applications in polynomial sampling. In particular, (40) implies that for any polynomial $p\in\mathscr{P}$ such that $\|p\|_{E}\leq 1$, one has
[…]
Example 7 (Chebyshev-Lobatto points as an admissible mesh).
Consider $[-1,1]$ as (the closure of) a Riemannian manifold isometric to the upper semicirle $\mathbb{S}^{1}_{+}:=\{x\in\mathbb{R}^{2}:|x|=1,x_{2}>0\}$ endowed by the round metric $g$ that, once pulled back onto the open interval $(-1,1)$, induces the distance $d_{E}(x,y):=|\operatorname*{arcos}(y)-\operatorname*{arcos}(x)|$. Assuming that $p(x)=\|p\|_{E}=1$ for some $x\in\operatorname*{int}E$, it is possible to rewrite (41) as
$|\operatorname*{arcos}p(y)|\leq\deg(p)d_{E}(x,y)\,,\quad\forall p\in\mathscr{P}.$
Hence, if $h=d_{E}(x,y)<\pi/(2\deg(p))$ we have $|p(y)|\geq\cos(\deg(p)h).$ Therefore, if $\mathcal{T}=\{x_{1},x_{2},\dots,x_{M}\}$ is a set of points in $E$ such that
$h(\mathcal{T},K):=\max_{y\in E}\min_{x\in\mathcal{T}}d_{E}(x,y)<\pi/(2r),$
then one has the sampling inequality
(42) $\max_{x\in E}|p(x)|\leq\frac{1}{\cos(rh(\mathcal{T},K))}\max_{x\in\mathcal{T}}|p(x)|$
for any $p\in\mathscr{P}_{r}.$ In particular a set $\mathcal{T}_{m}$ of $m+1$ Chebyshev-Lobatto points $x_{i}:=\cos(i\pi/(m+1))$ has the property $h(\mathcal{T}_{m},[-1,1])=\frac{\pi}{2m}$. It follows that $\max_{x\in[-1,1]}|p(x)|\leq\frac{1}{\cos(\pi r/(2m))}\max_{0\leq i\leq m}|p(x_{i})|$ for any $m=m(r)>r.$ In other words $\{\mathcal{T}_{m(r)}\}$ is an admissible mesh for $E.$
The generalization of the Bernstein Inequality (40) to the case of several variables has been carried out by Baran [6, 7]. The crucial step in this generalization consists in identifying the function $1/\sqrt{1-x^{2}}$ appearing in (40) as the normal derivative at $x$ of the Green function of $\mathbb{C}\setminus[-1,1]$ with a logarithmic pole at infinity. This quantity has a natural counterpart in several complex variables that, for a polynomially determining set $E\subset\mathbb{C}^{n}$, is the pluricomplex Green function
$V_{E}^{*}(z):=\limsup_{\zeta\to z}\sup\left\{\frac{1}{\deg p}\log|p(\zeta)|,p\in\mathscr{P},\,\max_{K}|p|\leq 1\right\},$
which is a maximal plurisubharmonic function solving the complex Monge Ampere homogeneous equation [48]. The generalization of (40) then assumes the following form.
Theorem 5 (Baran Inequality [6]).
Let $E\subset\mathbb{R}^{n}$ be a compact set with non empty interior. Then
(43) $\frac{|\partial_{v}p(x)|}{\sqrt{1-|p(x)|^{2}}}\leq\deg(p)\partial_{v}^{+}V_{E}^{*}(x),\quad\forall p\in\mathscr{P}:\,\max_{K}|p|\leq 1,\,v\in\mathbb{R}^{n},$
where
$\partial_{v}^{+}V_{E}^{*}(x):=\liminf_{t\to 0^{+}}V_{E}^{*}(x+itv).$
One can think of $F(x,v):=\partial_{v}^{+}V_{E}^{*}(x)$ as a metric on the tangent space to $\operatorname*{int}E$ at $x$. In general this is only a Finsler metric [15], but in few very relevant situations it turns out to be a Riemannian metric [51]. We term such a metric the Baran metric of $E$, and we denote it by $\delta_{E}(x,v)$. Via the Carnot-Caratheodory construction, $\delta_{E}(x,v)$ induces the distance
$d_{E}(x,y):=\inf\left\{\int_{0}^{1}\delta_{E}(\gamma(t),\gamma^{\prime}(t))dt,\gamma(0)=x,\ \gamma(1)=1,\ \operatorname*{supp}\gamma\subset K\right\},$
which we call Baran distance on $E$. Within this formalism, it is clear that equation (42) holds in the multivariate context, provided that necessary changes are made.
Theorem 6 (Fundamental estimate by Baran Inequality).
Let $E\subset\mathbb{R}^{n}$ be the closure of a bounded Lipshitz domain. Let $\mathcal{T}=\{T^{(s)}\}_{s=1,\dots,M}\subset\mathscr{A}^{k}(E)$ be such that, for any $s=1,2,\dots,M$, there exist $x^{(s)}\in E$, $A^{(s)}\in M_{n,k}(\mathbb{R}^{n})$ with $\operatorname*{Rank}(A^{(s)})=k$, a $\mathcal{L}^{k}$-measurable set $\Omega^{(s)}\subset\mathbb{R}^{k}$ with $0<\mathcal{L}^{k}(\Omega^{(s)})<+\infty$, for which
$T^{(s)}(\omega)=T_{S^{(s)}}(\omega)=\frac{1}{\mathcal{H}^{k}(S^{(s)})}\int_{S^{(s)}}\langle\omega;\sigma^{(s)}\rangle\mathrm{d}\mathcal{H}^{k}\;,\;\;\forall\omega\in\mathscr{D}_{0}^{k}(E)\,,$
where $S^{(s)}$ is the $k$-rectifiable set $\{x^{(s)}+A^{(s)}y,y\in\Omega^{(s)}\}\subset E$ endowed by the orientation
$\sigma^{(s)}:=\frac{A_{:,1}^{(s)}\wedge\dots\wedge A_{:,k}^{(s)}}{|A_{:,1}^{(s)}\wedge\dots\wedge A_{:,k}^{(s,r)}|}\,.$
Let $r\in\mathbb{N}$ and assume that there exists $c_{1}<1$ such that, for any $x\in E$, we can find $s_{1},s_{2},\dots,s_{m}\in\{1,\dots,N\}$ such that
(44) $c:=$ $\sup_{\tau\in\Lambda_{k}}\Bigg{\{}\min_{a\in\mathbb{R}^{m}}\Big{\{}|a|_{1}:\sum_{j=1}^{m}a_{j}\sigma^{(s_{j})}=\tau\Big{\}},|\tau|=1,\tau\text{ simple}\Bigg{\}}<+\infty$
(45) $h:=$ $\max_{j=1,\dots,m}\max_{z\in S^{s_{j}}}d_{E}(x,z)<\frac{\pi}{2r}\,.$
Then, for any $\omega\in\mathscr{P}_{r}\Lambda^{k}$ we have
(46) $\|\omega\|_{0}\leq\frac{c}{\cos(rh)}\|\omega\|_{\mathcal{T}}.$
Proof.
Let $\omega\in\mathscr{P}_{r}\Lambda^{k}$ be any polynomial form. We prove that (46) holds for $\omega$ under the assumption that there exist $\bar{x}\in E$ and a simple $\bar{\tau}\in\Lambda_{k}$ such that
(47) $\max_{x\in E}|\omega(x)|^{*}=\langle\omega(\bar{x});\bar{\tau}\rangle,$
the general case will easily follow by approximation.
Let $m\in\mathbb{N}$, $s_{1},s_{2},\dots,s_{m}\in\{1,\dots,N\}$ be such that
$\max_{j=1,\dots,m}\max_{z\in S^{s_{j}}}|\bar{x}-z|\leq h,$
and let $\bar{a}\in\mathbb{R}^{m}$ be any vector realizing
(48) $\min_{a\in\mathbb{R}^{m}}\Big{\{}|a|_{1}:\sum_{j=1}^{m}a_{j}\sigma^{(s_{j})}=\bar{\tau}\Big{\}}\leq c.$
Notice that $p(x):=\langle\omega(x);\bar{\tau}\rangle$ is a polynomial (whose degree does not exceed $r$) achiving its uniform norm on $E$ at the point $\bar{x}$. Exploiting (47) and (48) we compute
$\max_{x\in E}|\omega(x)|^{*}=$ $\langle\omega(\bar{x});\bar{\tau}\rangle=:p(\bar{x})\leq\frac{1}{\cos(rd_{E}(\bar{x},x))}p(x)\leq\frac{1}{\cos(rh)}\langle\omega(x);\bar{\tau}\rangle$
$=$ $\frac{1}{\cos(rh)}\sum_{j=1}^{m}\bar{a}_{j}\langle\omega(x);\sigma^{(s_{j})}\rangle\leq\frac{1}{\cos(rh)}\sum_{j=1}^{m}\bar{a}_{j}\frac{1}{\mathcal{H}^{k}(S^{s_{j}})}\int_{S^{s_{j}}}\langle\omega(x);\sigma^{(s_{j})}\rangle\mathrm{d}\mathcal{H}^{k}(x)$
$\leq$ $\frac{|\bar{a}|_{1}}{\cos(rh)}\|\omega\|_{\mathcal{T}}\,.$
This shows the claim. ∎
Example 8 (Admissible integral $k$-mesh for the $n$-cube by Baran Inequality).
Let $m>r$ and set $\tilde{x}_{i}:=\cos(i\pi/m)$ for any $i\in\{0,\dots,m\}.$ Consider the $n$-th cartesian product of $\{\tilde{x}_{0},\dots,\tilde{x}_{m}\}$ and the corresponding tessellation $\{Q_{j}\}_{j=1,\dots,m^{n}}$ of $E:=[-1,1]^{n}$ made of $n$-dimensional parallepipeds. Let $\mathcal{F}^{1,r},\dots,\mathcal{F}^{N,r}$, with $N=\binom{n}{k}m^{k}(m+1)^{n-k}$, be the collection of all the $k$-dimensional faces of the parallepipeds of the tessellation, with no repetitions. Any $k$-face $\mathcal{F}^{(s,r)}$ is of the form
$\mathcal{F}^{(s,r)}=\Big{\{}x\in E:x_{\alpha(j,s)}=\tilde{x}_{i(j,s)},\,j=1,\dots,n-k,\,\tilde{x}_{i(j,s)}\leq x_{\alpha(j,s)}\leq\tilde{x}_{i(j,s)+1},\,j=n-k+1,\dots,n\Big{\}}\,,$
where $\alpha(\cdot,s)$ is a permutation of the set $\{1,2,\dots,n\}$ and $i(j,s)\in\{0,\dots,m-1\}$ for any $j=1,2,n$, and $s=1,\dots,N.$ The tangent space to such a $k$-face is then spanned by $e_{\alpha(n-k+1,s)},\dots,e_{\alpha(n,s)}.$ Therefore we can give a parametrization and an orientation to $\mathcal{F}^{(s,r)}$ by setting, for each $j=1,\dots,n$ and $s=1,\dots,N$,
$x^{(s,r)}_{\alpha(j,s)}=$ $\tilde{x}_{i(j,s)}$
$A^{(s,r)}=$ [±$[\pm e_{\alpha(n-k+1,s)},\dots,\pm e_{\alpha(n,s)}]$
$\Omega^{(s,r)}=$ $\otimes_{j=n-k+1}^{n}[\tilde{x}_{i(j,s)},\tilde{x}_{i(j,s)+1}]\,,$
where signs are suitably chosen in order to have $\mathcal{F}^{(s,r)}=\{x^{(s,r)}+A^{(s,r)}y,\,y\in\Omega^{(s,r)}\}\subset E.$
Now, recalling that for the $n$ dimensional cube $E$ one has
$d_{E}(x,y):=\max_{i\in\{1,\dots,n\}}|\operatorname*{arcos}(x_{i})-\operatorname*{arcos}(y_{i})|,$
it is immediate to verify that $\mathcal{T}^{(r)}=\{(x^{(s,r)},A^{(s,r)},\Omega^{(s,r)}),s=1,\dots,N\}$ satisfies Eq. (44) with $c=\sqrt{\binom{n}{k}}$ and Eq. (45) with $h=\pi/(2m)$. Hence, due to Theorem 6, $\{\mathcal{T}^{(r)}\}$ is an admissible integral $k$-mesh of constant $C=\sqrt{\binom{n}{k}}/\cos(\frac{\pi r}{2m}).$
Remark 14.
In order to fairly compare the cardinality of the mesh constructed in Example 8 with that constructed in Example 6 (reported in Figure 2 for $c_{1}=1/2$), we need to set $h$ as the same constant of the mesh, i.e.,
$C=\frac{\sqrt{\binom{n}{k}}}{1-c_{1}}=\frac{\sqrt{\binom{n}{k}}}{\cos(\frac{\pi r}{2m})},$
that leads to $m=\lceil 3r/2\rceil$. For such a choice we have $M(r)\sim\binom{n}{k}\frac{3^{k}}{2^{n}}r^{k}(3r+2)^{n-k}$. The comparison of the cardinality of the two families, reported in in Figure 3 for the case $n=3$, $k=2$, shows that this latter strategy yields sensibly smaller sets.
Figure 3. Comparison of the cardinality of the mesh constructed in Example 8 with the one constructed in Example 6 in the case of $c_{1}=1/2$, $n=3$, and $k=2$. As a reference we depict also the curve $r\mapsto\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}.$
4.4. Admissible integral $k$-meshes for the $2$-simplex and the $3$-simplex by Baran Inequality
We dedicate this subsection to the explicit construction of a low cardinality admissible integral $k$-mesh for the $n$-simplex, with $n=2$ or $3$, relying upon Theorem 6. This algorithmic construction can be likewise extended to the case $n>3$, but requires a significantly heavier notation.
To this end, we consider $k$-faces of families of coordinate simplices, i.e. sets of the form
$S(\tau,a,\ell):=\Big{\{}x\in\mathbb{R}^{n}:\,\tau_{i}(x_{i}-a_{i})\geq 0,\,\sum_{i=1}^{n}\tau_{i}(x_{i}-a_{i})\leq\ell\Big{\}},$
and propose the following notation for the $k$-integral meshes we construct.
Definition 4.
Let $\tilde{\mathcal{T}}_{k,m,n}$ denote
  • For $\boldsymbol{n=2}$
    • for $k=1$, the collection of edges of simplices $S(\tau,a,\ell)$, with $\tau=\{1,1\}$, $\ell=1/m$, and $a\in((m-1)\Sigma_{n}\cap\mathbb{N}^{2})=\{(i,j)\in\mathbb{N}^{2}:i+j\leq m-1\}$
    • for $k=2$, the collection of $2$-simplices of the form $S(\tau,a,\ell)$, with either $\tau=\{1,1\}$, $\ell=1/m$, and $a/\ell\in((m-1)\Sigma_{n}\cap\mathbb{N}^{2}),$ or $\tau=\{-1,-1\}$, $\ell=1/m$, and $(a+\tau)/\ell\in(m-1)\Sigma_{n}\cap\mathbb{N}^{2}.$
  • For $\boldsymbol{n=3}$
[…]
    • for $k=n$ the collection of the $n$-simplices introduced in the case $k<n$ and the collection of the tetrahedrons $T(a,\ell)$ of verices $a+(\ell,0,0)$, $a+(0,\ell,0)$, $a+(0,0,\ell)$,$a+(\ell,\ell,\ell)$, with $\ell=1/m$ and $a/\ell\in(m-1)\Sigma_{n}\cap\mathbb{N}^{3}$.
Figure 4 illustrates the elements of the tessellation proposed in Definition 4 for the case $n=3$.
Figure 4. Elements used in Definition 4 in the case $n=3$. Left: the four type of simplices that are always (i.e., $k=1,2,3$) considered. Right: the tetrahedron that is taken into account only in the case $k=3$.
Refer to caption Refer to caption
As shown in [6, 15], the Baran distance on the standard $n$-dimensional simplex $\Sigma_{n}:=\{x\in\mathbb{R}^{n}:x_{i}\geq 0,\,\sum_{i=1}^{n}x_{i}\leq 1\}$ has the explicit form
$d_{\Sigma_{n}}(x,y)=2\operatorname*{arcos}\left(\phi(x),\phi(y)\right),$
where $\phi:\Sigma_{n}\rightarrow\mathbb{S}_{n}\cap\mathbb{R}_{+}^{n+1}$ is defined as
$\phi(x_{1},\dots,x_{n})=\left(\sqrt{1-\sum_{i=1}^{n}x_{i}},\sqrt{x_{1}},\dots,\sqrt{x_{n}}\right)$
and $d_{\Sigma_{n}}$ is the Carnot-Caratheodory distance induced by the Riemannian metric
$g_{\Sigma_{n}}(x):=\left[\begin{array}[]{ccccc}x_{1}^{-1}&0&\dots&\dots&0\\ 0&x_{2}^{-1}&0&\dots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\dots&\dots&0&x_{n}^{-1}\end{array}\right]+\frac{1}{1-\sum_{i=1}^{n}x_{i}}\left[\begin{array}[]{ccccc}1&1&\dots&\dots&1\\ 1&1&\dots&\dots&1\\ 1&1&\dots&\dots&1\\ 1&1&\dots&\dots&1\\ 1&1&\dots&\dots&1\end{array}\right].$
defined on the interior of $\Sigma_{n}.$
Remark 15.
The volume form $d\operatorname*{Vol}_{\Sigma_{n}}=\frac{1}{\sqrt{(1-\sum_{i=1}^{n}x_{i})\prod_{i=1}^{n}x_{i}}}\mathrm{d}\mathcal{H}^{n}(x)$ induced by $g_{\Sigma_{n}}$ blows up around the boundary of $\Sigma_{n}$. As a consequence, two points with fixed Euclidean distance present a large Baran distance as they approach the boundary of $\Sigma_{n}$.
Since we aim at constructing an integral $k$-mesh with maximal diameter of the elements of the tessellation bounded above by $h>0$, this justifies the following adaptive strategy for the development of Algorithm 1:
  • We compute a suitable $m$ such that the maximal diameter of elements of the tessellation defining $\tilde{\mathcal{T}}_{k,m,n}$ is bounded above by $h$;
  • We add to our integral $k$-mesh all the $k$-faces of elements $E_{\beta}$ of this tessellation such that $E_{\beta}\cap\partial\Sigma_{n}\neq\emptyset$, i.e., elements with larger Baran diameter;
  • We iteratively repeat this procedure on the set until $\Sigma_{n}$ has been exhausted. Notice that in fact $\Sigma_{n}\setminus\cup_{\beta}E_{\beta}$ is still a simplex.
Such a procedure is formalized in Algorithm 1, and graphically depticted in Figure 5.
Algorithm 1 Simplex $k$-Mesh
1:$n\in\mathbb{N},$ $k\in\{0,n\}$, $r\in\mathbb{N}$, $h\in]0,\pi/(2r)[$
2:$c=0$, $s=1$, $T=\emptyset$, $p=1-\cos(h/2)$
3:while $s>0$ do
4:     $m\leftarrow\left\lceil\frac{s}{p\left(2\sqrt{c/p}+1\right)}\right\rceil$
5:     $\ell\leftarrow s/m$
6:     $\tilde{T}\leftarrow c+s\tilde{T}_{k,m,n}$
7:     $c\leftarrow c+\ell$
8:     $s\leftarrow s-(n+1)\ell$
9:     for $E\in\tilde{T}$ do
10:         if $E\not\subseteq c+s\Sigma_{n}$ then
11:              for $F\in\{k\text{-faces of }E\}$ do
12:                  $T\leftarrow T\cup F$
13:              end for
14:         end if
15:     end for
16:end while
17:$T$
Proposition 4.
Let $\mathcal{T}^{(r)}$ be the sequence of outputs of Algorithm 1 executed for each $r\in\mathbb{N}$ with $h:=\theta\pi/(2r)$ with $10<\theta<$ 1. Then $\mathcal{T}^{(r)}$ is an admissible integral $k$-mesh for the simplex with constant $C:=\frac{\sqrt{\binom{n}{k}}}{\cos(\theta\pi/2)}.$
Proposition 4 ensures that Algorithm 1 produces the desired admissible integral $k$-mesh. We present the steps of the proof as separate lemmata, and postpone the rest of the proof to the end of the section.
Figure 5. The costruction carried out by Algorithm 1. At each step the elements lyining in the grey area are not taken into account, in the next step such region is re-meshed by simplices having sides of updated length.
Refer to caption Refer to caption Refer to caption
The first step we need to accomplish in order to prove Proposition 4 is to compute, for any $k$ and $r$, a suitable $m$ for which we can apply Theorem 6 in which estimates (44) and (45) are written for the averaging currents defined by the $k$-faces of the element of the considered tessellation $\mathcal{T}_{k,m,n}$ that contains the test point $x\in\Sigma_{n}$. The following property of the Baran distance on the simplex will play a pivotal role in our construction.
Lemma 5.
Let $f:\Sigma_{n}\times\Sigma_{n}\rightarrow R$ be defined as $f(x,y)=\cos(d_{\Sigma_{n}}(x,y)/2)$. Then the functions $f(\cdot,y)$ and $f(x,\cdot)$ are concave on $\Sigma_{n}$. In particular the Baran diameter of any convex polytope $\mathcal{P}\subset\Sigma_{n}$ is the maximum Baran distance of its vertices.
Proof.
We show that $x\mapsto f(x,y)$ is convcave and the concavity of $y\mapsto f(x,y)$ will follow by simmetry. Indeed we can compute
$\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}(x,y)=-\frac{1}{4}\left(\delta_{i,j}\sqrt{\frac{y_{i}}{x_{i}^{3}}}+\frac{\sqrt{1-\sum_{h=1}^{n}y_{h}}}{\sqrt{1-\sum_{h=1}^{n}x_{h}}}\right).$
Therefore
$\sum_{i,j=1}^{n}v_{i}v_{j}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}(x,y)=$ $-\frac{1}{4}\left(\sum_{i=1}^{n}\sqrt{\frac{y_{i}}{x_{i}^{3}}}v_{i}^{2}+\frac{\sqrt{1-\sum_{h=1}^{n}y_{h}}}{\sqrt{1-\sum_{h=1}^{n}x_{h}}}\sum_{i=1}^{n}\sum_{j=1}^{n}v_{i}v_{j}\right)$
$=$ $-\frac{1}{4}\left(\sum_{i=1}^{n}\sqrt{\frac{y_{i}}{x_{i}^{3}}}v_{i}^{2}+\frac{\sqrt{1-\sum_{h=1}^{n}y_{h}}}{\sqrt{1-\sum_{h=1}^{n}x_{h}}}(\sum_{i=1}^{n}v_{i})^{2}\right)\leq 0.$
Let us denote by $\operatorname*{diam}_{\Sigma_{n}}$ the diameter with respect to the Baran distance. Notice that
$\frac{\operatorname*{diam}_{\Sigma_{n}}(\mathcal{P})}{2}=$ $\max_{x\in\mathcal{P}}\max_{y\in\mathcal{P}}\operatorname*{arcos}\left(\phi(x),\phi(y)\right)=\max_{x\in\mathcal{P}}\operatorname*{arcos}\min_{y\in\mathcal{P}}\left(\phi(x),\phi(y)\right)=\operatorname*{arcos}\min_{x\in\mathcal{P}}\min_{y\in\mathcal{P}}\left(\phi(x),\phi(y)\right)$
$=$ $\operatorname*{arcos}\min_{x\in\mathcal{P}}\min_{y\in\mathcal{P}}f(x,y)=\operatorname*{arcos}\min_{x\in\mathcal{P}}\min_{y\in\operatorname*{Extr}\mathcal{P}}f(x,y)=\operatorname*{arcos}\min_{x\in\operatorname*{Extr}\mathcal{P}}\min_{y\in\operatorname*{Extr}\mathcal{P}}f(x,y),$
where we denoted by $\operatorname*{Extr}\mathcal{P}$ the set of extremal points (hence vertices) of the convex polytope $\mathcal{P}$, which is indeed the set of vertices. Note that the last two equalities follow by the above proven concavity of $x\mapsto f(x,y)$ and from the concavity of $y\mapsto\min_{x\in\operatorname*{Extr}\mathcal{P}}f(x,y).$
We then give an asymptotic upper bound to the Baran diameter of the elements of a tessellation.
Lemma 6.
Let $n=2$ or $n=3$ and let $\mathcal{A}$ be the set of the elements of the tessellation $\tilde{\mathcal{T}}(k,m,n)$ described in Definition 4. Then
(49) $\max_{A\in\mathcal{A}}{\operatorname*{diam}}_{\Sigma_{n}}(A)=2\operatorname*{arcos}(1-1/m)\,.$
Proof.
First, notice that, for the case of tessellations containing also tetrahedrons, we do not need to take them into account. Indeed, due to Lemma 5, if $\mathcal{P}$ is a tetrahedron, its diameter is the distance of its most separated vertices, but, for any pair $(v_{1},v_{2})$ of such vertices, there exists a simplex $S$ of the same tessellation such that $u$ and $v$ are vertices of $S$. Hence, removing $\mathcal{P}$ from the maximization in (49) does not change the value of the maximum diameter.
Now, consider the elements $A\in\mathcal{A}(\tau)$ being simplices with the same $\tau$. Then we have
$\max_{A\in\mathcal{A}(\tau)}{\operatorname*{diam}}_{\Sigma_{n}}(A)=\max_{ma\in(m-1)\Sigma_{n}\cap\mathbb{N}^{n}}{\operatorname*{diam}}_{\Sigma_{n}}S(\tau,a,1/m)$
$=$ $2\operatorname*{arcos}\big{(}\min_{ma\in(m-1)\Sigma_{n}\cap\mathbb{N}^{n}}\min_{x,y\in S(\tau,a,1/m)}f(x,y)\big{)}$
$=$ $2\max_{a\in\operatorname*{Extr}[((m-1)\Sigma_{n}\cap\mathbb{N}^{n})/m]}\operatorname*{arcos}\big{(}\min_{x,y\in S(\tau,a,1/m)}f(x,y)\big{)}$
$=$ $2\max_{a\in\operatorname*{Extr}[((m-1)\Sigma_{n}\cap\mathbb{N}^{n})/m]}\operatorname*{arcos}\big{(}\min_{x,y\in\operatorname*{Extr}S(\tau,a,1/m)}f(x,y)\big{)},$
where in the last two lines we used the fact that $a\mapsto\min_{x,y\in S(\tau,a,1/m)}f(x,y)=\min_{x,y\in S(\tau,0,1/m)}f(x+a,y+a)$ is concave, which is a straightforward consequence of Lemma 5.
Then equation (49) is obtained by direct computation for each $\tau$ and using the simmetry of the problem. In order to clarify the procedure we report only the computation for the case $\tau=(1,\dots,1)$, $a=0$, since the other cases are completely analogous. In such a case one has
$\min_{x,y\in\operatorname*{Extr}S(\{1,\dots,\},0,1/m)}f(x,y)=\min\left\{\min\{f(0,e_{i}/m)\},\min\{f(e_{i}/m,e_{j}/m),i\neq j\}\right\}$
$=$ $\min\left\{\left\|\begin{pmatrix}\sqrt{1-1/m}\\ 0\\ \vdots\\ 0\\ \sqrt{1/m}\\ 0\\ \vdots\end{pmatrix}\right\|^{2}\,,\,\left(\begin{pmatrix}\sqrt{1-1/m}\\ 0\\ \vdots\\ \sqrt{1/m}\\ 0\\ 0\\ \vdots\end{pmatrix},\begin{pmatrix}\sqrt{1-1/m}\\ 0\\ \vdots\\ 0\\ \sqrt{1/m}\\ 0\\ \vdots\end{pmatrix}\right)\right\}$
$=$ $\min\{1,(1-1/m)\}=1-1/m\,,$
and the claim is proved. ∎
Completely analogous computations can be used to bound the maximum Baran diameter of the elements of the similarly defined tessellation of certain subsimplices, giving the following result.
Corollary 3.
Let $0<c<1/(n+1)$, $n=2$ or $n=3$, and let $\mathcal{A}$ be the set of the elements of the tessellation $(1-c(n+1))\tilde{\mathcal{T}}(k,m,n)$ of $(1-c(n+1))\Sigma_{n}$, with $\tilde{\mathcal{T}}(k,m,n)$ described in Definition 4. Then
(50) $\max_{A\in\mathcal{A}}{\operatorname*{diam}}_{\Sigma_{n}}(A)=2\operatorname*{arcos}\left(1-\left(\sqrt{c+\frac{s}{m}}-\sqrt{c}\right)^{2}\right)\,,$
where $s=1-(n+1)c.$
of Proposition 4.
Pick any $r\in\mathbb{N}$ and $h<\frac{\pi}{2r}$. It is clear by construction and by Eq. (49) and Eq. (50) that Algorithm 1 will provide an integral $k$-mesh $\mathcal{T}^{(r)}$ such that for any $x\in\Sigma_{n}$, we can find either simplex or a tetrahedron $S$ with $\operatorname*{diam}_{\Sigma_{n}}(S)\leq h$, such that all its $k$-faces are in $\mathcal{T}^{(r)}$, and $x\in S.$ Hence the condition (45) of Theorem 6 is satisfied.
If $S$ is a coordinate simplex, then it is clear that (44) of Theorem 6 is satisfied as well with $c=\sqrt{\binom{n}{k}}.$ If $S$ is one of the tetrahedrons described in Definition 4, hence we are considering the case $k=n$, we can take $c=1$, since in such a case $\Lambda^{k}\cong\mathbb{R}$, all $k$-vectors are in particular simple and the integer $m$ in (44) is $1$.
Now we need to provide an asymptotic upper bound for the cardinality of $\mathcal{T}^{(r)}$ as $r\to+\infty$. Notice that the sequence of real numbers $\ell$ computed by Algorithm 1 is non decreasing. At the first iteration, Algorithm 1 starts from a tessellation of the simplex made of simplices of side $\ell_{1}=1/\lceil 1/p\rceil$, hence it considers
$M_{1}=\frac{\mathcal{L}^{n}(\Sigma_{n})}{\mathcal{L}^{n}(\ell_{1}\Sigma_{n})}=\ell_{1}^{-n}=\lceil 1/p\rceil^{n}$
simplices. Then, at each iteration the sub-tessellation covering the simplex $(1-(n+1)\sum_{l=1}^{j-1}\ell_{j})\Sigma_{n}$ is replaced by a tessellation made by simplices of larger side, i.e., the considered tessellation has smaller cardinality. Therefore we have
$\operatorname*{Card}\mathcal{T}^{(r)}\leq\binom{n+1}{k}M_{1}=\binom{n+1}{k}\lceil 1/p\rceil^{n},$
where $\binom{n+1}{k}$ is the number of $k$-faces of an $n$ simplex. One finally has
$\lceil 1/p\rceil^{n}\sim\left(\frac{1}{1-\cos\frac{h}{2}}\right)^{n}=\left(\frac{1+\sqrt{\frac{1+\cos h}{2}}}{1-\frac{1}{2}-\frac{\cos h}{2}}\right)^{n}\sim\left(\frac{2}{h^{2}}\right)^{n}=\left(\frac{2}{\theta^{2}\pi^{2}}r^{2}\right),$
as $r\to+\infty.$ Since the set $\mathcal{T}^{(r)}\subset\mathscr{A}^{k}(\Sigma_{n})$ satisfies conditions (44) and (45) of Theorem 6, and has a cardinality that grows polynomially with respect to $r$, the sequence $\{\mathcal{T}^{(r)}\}$ is an admissible integral $k$-mesh for $\Sigma_{n}.$
Remark 16.
We stress that the asymptotic bound for the cardinality of the output of Algorithm 1 provided in the proof of Proposition 4 is extremely rough. The corresponding integral $k$-mesh in fact presents a cardinality which increases proportionally to $\log(r)\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}$, as depicted in Fig. 6. In the nodal case $k=0$, meshes satisfying such a cardinality growth rate are termed quasi-optimal, while meshes with the cardinality growing as a multiple of $\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}$ are termed optimal, see, e.g., [27, 44, 50].
Figure 6. Left: cardinalities $\operatorname*{Card}\mathcal{T}^{(r)}$ of the admissible integral $k$-mesh constructed by Algorithm 1 (with $n=2$, $k=1$, $\theta=2/3$, so that $h=\theta\pi/(2r)$ has the property $21/\cos(hr)=$ 2) compared with the growth of $r^{2}\log(r)$. Right: the ratio $(\operatorname*{Card}\mathcal{T}^{(r)})/(r^{2}\log(r)).$
Refer to caption Refer to caption
5. Interpolation and fitting on admissible integral $k$-meshes: error estimates revisited
5.1. Extremal sets of currents of Fekete and Leja type
It is clear from Section 2 that a good interpolation scheme needs to have slowly increasing Lebesgue constant. Already in the case $k=0$ and $n=1$ (i.e., nodal interpolation of univariate functions), the minimization of the Lebesgue constant among all possible set of interpolation nodes on a given compact set $E$ (the Lebesgue problem) becomes unmanageable already for mild values of $r$. The quest for computable sub-optimal solutions is thus of major interest in approximation theory.
Fekete points are the most studied sub-optimal solution of the Lebesgue problem. A set of nodes $X:=\{x_{1},x_{2},\dots,x_{N}\}\subset E^{N}$ with $N:=\operatorname*{dim}\mathscr{P}_{r}$, is termed a set of Fekete points for $E$ if it maximizes the determinant of the Vandermonde matrix $\operatorname*{vdm}$:
(51) $|\det\operatorname*{vdm}(X,\mathcal{B}_{r})|=\max_{Y\in E^{N}}|\det\operatorname*{vdm}(Y,\mathcal{B}_{r})|,\;\;\text{with }\operatorname*{vdm}(Y,\mathcal{B}_{r})_{i,j}=b_{j}(y_{i}),$
for one, and thus for all, basis $\mathcal{B}_{r}=\{b_{1},\dots,b_{N}\}$ of $\mathscr{P}_{r}.$ Fekete problems in the segmental framework have also been studied [21].
Another proposed sub-optimal solution to the Lebesgue problem are Leja sequences. Leja sequences are defined by means of the greedy version of the maximization procedure appearing in (51). Precisely one first picks $x_{1}\in E$, then iteratively sets, for $k\in\mathbb{N}$,
(52) $x_{k+1}\in{\operatorname*{argmax}}_{x\in E}|\det\operatorname*{vdm}(\{x_{1},\dots,x_{k},x\},\{b_{1},\dots,b_{k+1}\})|\,.$
The first interest on Fekete points is easy to see: if $\{x_{1},x_{2},\dots,x_{N}\}$ are Fekete points, then we can write
$\mathscr{L}(X,\mathscr{P}_{r})=\max_{x\in E}\sum_{i=1}^{N}|\ell_{i}(x)|=\max_{x\in E}\sum_{i=1}^{N}\frac{|\det\operatorname*{vdm}(X^{(i)}(x),\mathcal{B}_{r})|}{|\det\operatorname*{vdm}(X,\mathcal{B}_{r})|}\leq\sum_{i=1}^{N}1=N,$
where we denoted by $\ell_{i}$ the $i$-th Lagrange polynomial and by $X^{(i)}(x)$ the set $X$ with the $i$-th element replaced by $x$. Hence Fekete points have a polynomially increasing Lebesgue constant.
The relationship between Lebesgue problem and the maximization of the Vandermonde determinant also bridges approximation theory with logarithmic potential theory when $n=1$ [59, 57] and pluripotential theory when $n>1$ [48]. This link is offered by the asymptotic of the optimal Vandermonde determinant as $r\to+\infty$, which is the transfinite diameter of the set $E$:
(53) $\delta(E)=\lim_{r}\delta^{(r)}(E):=\left[\max_{\boldsymbol{x}\in E^{N}}\Big{|}\operatorname*{vdm}(\{x_{1},\dots,x_{N}\},\mathcal{B}_{r}^{\text{mon}})\Big{|}\right]^{1/(l_{r})},$
where $l_{r}:=\sum_{j=1}^{r}[j(\operatorname*{dim}\mathscr{P}_{j}-\operatorname*{dim}\mathscr{P}_{j-1})],$ and $\mathcal{B}_{r}^{\text{mon}}$ is the monomial basis introduced in Section 2.3. The existence of the limit in the definition of $\delta(E)$ is rather straightforward if $n=1$, while was proved for $n>1$ by Zaharjuta [65] with more complicated technologies. In one complex variable, $\delta(E)$ is the logarithmic capacity of $E$, the quantity that distinguishes (from the potential-theoretic point of view) relevant sets from negligible ones. Arrays of interpolation nodes leading to the transfinite diameter of $E$ are termed asymptotically Fekete and necessarily tend, in the weak topology of measures, to $\mu_{E}$, the equilibrium measure of the compact set $E$.
When $n>1$ the situation becomes more involved. Nevertheless, all the above mentioned relations among Fekete points, transfinite diameter, and suitably generalized versions of equilibrium measure, Green function, and capacity, have been extended replacing the potential-theoretic point of view by the framework of pluripotential theory. In particular, the sequence of uniform probability measures supported at Fekete points converges to $\mu_{E}$, the pluripotential equilibrium measure of $E$ [8, 9].
A further step is required when the geometry is enriched and differential forms are considered. In [22] authors introduced the (weighted) transfinite diameter of $E$ with respect to a real vector space $U$. For that, a choice of an orthonormal basis of $U$ is required. This definition can be specialized to the case of $U=\Lambda^{k}$ via the orthonormality introduced in Eq. (12), leading to the following definition:
$\delta(E,\Lambda^{k}):=$ $\lim_{r}\delta^{(r)}(E,\Lambda^{k})$
(54) $\delta^{(r)}(E,\Lambda^{k}):=$ $\max_{\mathcal{T}\in[\mathscr{S}_{0}^{k}(E)]^{N(r)}}|\det\operatorname*{vdm}(\mathcal{T},\mathcal{B}_{r}^{\text{mon},k})|^{1/(\binom{n}{k}l_{r})},$
where $\mathscr{S}_{0}^{k}(E)$ is the unit sphere of $\mathscr{M}_{0}^{k}(E)$ defined in (4), and $\mathcal{B}_{r}^{\text{mon},k}$ is the monomial basis of polynomial forms of degree at most $r$ and order $k$, i.e., $\{x^{\beta}dx_{\alpha},\,|\beta|\leq r,\,|\alpha|=k\}.$ The existence of the limit in equation (54) is proved in [22] in a weighted setting. Further, the weighted transfinite diameter with respect to $\Lambda^{k}$ can be expressed as a geometric mean of standard weighted transfinite diameters, simplifying in the unweighted case to
$\delta(E,\Lambda^{k})=\delta(E).$
With this at hand, we can extend the definition of Fekete points and asymptotically Fekete arrays.
Definition 5 (Fekete currents).
Let $\mathcal{F}^{(r)}:=\{\mathcal{F}^{(1,r)},\dots,\mathcal{F}^{(N(r),r)}\}\subset\mathscr{S}_{0}^{k}(E)$, $N(r)=\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}$. Then $\{\mathcal{F}^{(1,r)},\dots,\mathcal{F}^{(N(r),r)}\}$ are termed Fekete currents if
$|\det\operatorname*{vdm}(\mathcal{F}^{(r)},\mathcal{B}_{r}^{\text{mom},k})|=\max_{\mathcal{S}\in[\mathscr{S}_{0}^{k}(E)]^{N(r)}}|\det\operatorname*{vdm}(\mathcal{S},\mathcal{B}_{r}^{\text{mon},k})|.$
The sequence {~$\{\widetilde{\mathcal{F}}^{(r)}\}_{r\in\mathbb{N}}$ is termed asymptotically Fekete if
(55) $\lim_{r}|\det\operatorname*{vdm}(\widetilde{\mathcal{F}}^{(r)},\mathcal{B}_{r}^{\text{mom},k})|^{1/\left(\binom{n}{k}l_{r}\right)}=\delta(E).$
Since the asymptotics discussed above also hold for asymptotically Fekete arrays, see [11], Leja sequences defined in (52) can be similarly extended to the currents based setting. As for the case $k=0$, this definition depends on the ordering of the polynomial basis.
Definition 6 (Leja sequences of currents).
Let $\mathcal{Q}=\{q_{1},q_{2},\dots\}$ be a lower triangular basis of $\mathscr{P}\Lambda^{k}=\cup_{r\in\mathbb{N}}\mathscr{P}_{r}\Lambda^{k}$ in the sense of Definition 1, and, for any $N\in\mathbb{N}$, denote by $\mathscr{V}_{N}$ the linear space $\operatorname*{span}\{q_{1},\dots,q_{N}\}$. Let $T_{1}\in\operatorname*{argmax}\{|T(q_{1})|,\,T\in\mathscr{S}_{0}^{k}(E)\}$, and let, for any $i\in\mathbb{N}$, $T_{i+1}\in\mathscr{S}_{0}^{k}(E)$ be chosen such that
$L_{i+1}\in\operatorname*{argmax}\{|\det\operatorname*{vdm}(\{L_{1},\dots,L_{i},L\},\mathscr{V}_{i+1})|,\,L\in\mathscr{S}_{0}^{k}(E)\}\,.$
The sequence $\{L_{i}\}$ is a Leja sequence for $\mathscr{P}\Lambda^{k}$ relative to $\mathcal{Q}$.
The problem of finding Fekete currents has some simplifications with respect to the Lebesgue problem. Nevertheless, it is still completely unfeasible from a computational point of view. Polynomial admissible mesh have been successifully employed to obtain a further simplification in the case of nodal interpolation of functions [12]. Indeed, one may replace, in Definition 5 and Definition 6, the domain of the maximization $\mathscr{S}_{0}^{k}$ with any admissible integral $k$-mesh. In such a case, we will refer to the corresponding currents as Fekete and Leja sequences extracted from the mesh. Also in the context of the present work this approach is particularly profitable, as shown by the following theorem.
Theorem 7.
Let $E\subset\mathbb{R}^{n}$ be a compact non-pluripolar set, $k\in\{1,\dots,n\}$ and let $\{\mathcal{T}^{(r)}\}$ be an admissible integral $k$-mesh for $E$ of constant $C$. Any sequence of Fekete or Leja currents extracted from the mesh $\{\mathcal{T}^{(r)}\}$ is asymptotically Fekete. Moreover, when {~$\{\widetilde{\mathcal{F}}^{(r)}\}$ is a Fekete sequence extracted from the mesh,
(56) (~$\mathscr{L}(\widetilde{\mathcal{F}}^{(r)},\mathscr{P}_{r}\Lambda^{k})\leq C\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}.$
In particular ((~$\left(\mathscr{L}(\widetilde{\mathcal{F}}^{(r)},\mathscr{P}_{r}\Lambda^{k})\right)^{1/r}\leq\left(C\binom{n}{k}\binom{n+r}{r}\right)^{1/r}\sim\left(\frac{C}{(n-k)!k!}r^{n}\right)^{1/r}\to 1$ as $r\to+\infty$.
Proof.
Let $\{\mathcal{F}^{(r)}\}$ be any true Fekete sequence of currents for $E$. Let {~$\{\widetilde{\mathcal{F}}^{(r)}\}$ be the Fekete currents extracted from the mesh $\mathcal{T}^{(r)}$ and $\{\widetilde{L}_{i}\}$ be the Leja sequence extracted from the same mesh. Using Definition 5 and Definition 6 and the sampling property (29), one can write
$|\det\operatorname*{vdm}({\mathcal{F}}^{(r)},\mathcal{B}_{r}^{\text{mom},k})|$ $\geq|\det\operatorname*{vdm}(\widetilde{\mathcal{F}}^{(r)},\mathcal{B}_{r}^{\text{mom},k})|\geq|\det\operatorname*{vdm}(\{\widetilde{L}_{1},\ldots,\widetilde{L}_{N(r)}\},\mathcal{B}_{r}^{\text{mom},k})|$
$\geq\frac{1}{C^{N(r)}}|\det\operatorname*{vdm}({\mathcal{F}}^{(r)},\mathcal{B}_{r}^{\text{mom},k})|.$
Taking the $\binom{n}{k}l_{r}$-root of each term, one sees that the first and the last terms satisfy (55), and so do all the intermediate terms.
To prove the second part of the statement, consider the Lagrange basis $\{\omega_{1},\ldots,\omega_{N(r)}\}$ associated with the set of currents ~$\widetilde{\mathcal{F}}^{(r)}$. It follows from the definition of Fekete currents extracted from $\mathcal{T}^{(r)}$ that
$\|\omega_{i}\|_{\mathcal{T}^{(r)}}=\max_{s}\frac{\left|\det\operatorname*{vdm}(\{\widetilde{F}^{(1,r)},\ldots,\widetilde{F}^{(i-1,r)},T^{(s,r)},\widetilde{F}^{(i+1,r)},\ldots,\widetilde{F}^{(N,r)}\},\mathcal{B}_{r}^{\text{mom},k})\right|}{|\det\operatorname*{vdm}({\widetilde{\mathcal{F}}}^{(r)},\mathcal{B}_{r}^{\text{mom},k})|}\leq 1.$
Using the sampling property (29) one gets $\|\omega_{i}\|_{0}\leq C\|\omega_{i}\|_{\mathcal{T}^{(r)}}$, whence taking the sum for $i=1,\ldots,N(r)$, the claim follows. ∎
Remark 17.
To the best of the authors’ knowledge, Theorem 7 provides the first example of interpolation schemes based on integration of differential forms having Lebesgue constant of polynomial growth. In particular, the classical Lebesgue estimate for interpolation, which has been discussed in Section 3, in the setting of Theorem 7 reads
$\|\omega-\Pi^{(r)}\omega\|_{0}\leq\left(1+C\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}\right)d_{r}(\omega,E),$
where $d_{r}(\omega,E):=\in\{\|\omega-\theta\|_{0},\,\theta\in\mathscr{P}_{r}\Lambda^{k}\}$ is the error of best approximation, and $\Pi^{(r)}$ is the interpolation operator introduced in Eq. (11), based on the supports provided by Theorem 7.
The search of Fekete (or Leja) currents extracted from an integral admissible $k$-mesh instead of true Fekete (or Leja) currents suggested by Theorem 7 moves the problem from the context of continuous optimization to the discrete one. However, the problem on the discrete level is still NP-hard [26]. For this reason, heuristic or stochastic approaches are generally pursued [13, 12]. All the new objects introduced in the present paper have been suitably defined in order to extend to differential forms the constructions of approximate Fekete points (AFP algorithm) or of discrete Leja points (DLP algorithm) proposed in [13].
Remark 18.
Consistetly with the nodal case, the AFP and DLP algorithms provably extract collections of asymptotically Fekete currents for differential forms. This is easily achieved reproducing the proof of [12, Thm. 1]. In contrast, even for $k=0$, it has not been proven that the algorithms output sequences of currents having sub-exponentially growing Lebesgue constant, i.e., $\limsup_{r\to\infty}\left(\mathscr{L}(\widetilde{\mathcal{F}}^{(r)},\mathscr{P}_{r}\Lambda^{k})\right)^{1/r}\leq 1$. Nevertheless, numerical experiment carried in the nodal framework suggest that such Lebesgue constants exhibit polynomial growth [14].
An example of Fekete currents extracted by the AFP algorithm is depicted in Figure 7.
Figure 7. Two examples of approximate Fekete currents extracted by the AFP algorithm from an admissible integral $1$-mesh on the unit $2$-simplex. Left: $r=5$, $C=3$, $\mathscr{L}\approx 3.8$. Right: $r=7$, $C=3$, $\mathscr{L}\approx 7$.
Refer to caption Refer to caption
5.2. A remark on the asymptotics of Fekete currents
Given a sequence of true Fekete points $\{x_{1}^{(r)},\dots,x_{N(r)}^{(r)}\}$, the results of [8, 9] imply the asymptotic of empirical measures
$\mu^{(r)}:=N(r)^{-1}\sum_{i=1}^{N(r)}\delta_{x_{i}^{(r)}}\rightharpoonup\mu_{E}\;\;\text{ as }r\to+\infty\;.$
The main result of [12, Thm. 1] extends the above asymptotics to approximate Fekete points and discrete Leja sequences. The analogous property is not contained in the claim of Theorem 7, when treating the case $k\geq 1$. This discrepancy depends on the lack of a strong asymptotic behavior of Fekete currents. This gap has been only partially filled by the extension of the deep results of [8, 9] to polynomial differential forms, which has been carried out in [22].
Indeed, in [22, Cor. 5.1] it is shown that, if Fekete currents $\{\mathcal{F}^{(r)}\}=\{\mathcal{F}^{1,r},\dots,\mathcal{F}^{N(r),r}\}$ are of the form $\mathcal{F}^{s,r}:=\sigma_{i(s)}\delta_{x_{s}^{(r)}}$ for some orthonormal basis $\{\sigma_{1},\dots,\sigma_{\binom{n}{k}}\}$ of $\Lambda_{k}$, then the empirical current $T^{(r)}:=\frac{1}{N(r)}\sum_{s=1}^{N(r)}\mathcal{F}^{s,r}$ associated to $\mathcal{F}^{(r)}$ satisfy
$T^{(r)}\rightharpoonup\frac{1}{\binom{n}{k}}\sum_{i=1}^{\binom{n}{k}}\sigma_{i}\,\mu_{E}\,,$
[…]
One of the difficulties in proving an asymptotic behaviour of asymptotically Fekete currents is due to the higher geometric richness of the context of currents with respect to the one of Borel measures. We hence conjecture the following:
Conjecture 1.
Let $\{\mathcal{F}^{(r)}\}=\{\mathcal{F}^{1,r},\dots,\mathcal{F}^{N(r),r}\}$, $\mathcal{F}^{s,r}\in\mathscr{A}^{k}(E)$, for any $s=1,\dots,N(r)$, be an asymptotically sequence of currents for the compact, non-pluripolar set $E\subset\mathbb{R}^{n}$. Denote by
$T^{(r)}:=\frac{1}{N(r)}\sum_{s=1}^{N(r)}\mathcal{F}^{s,r}$
the empirical current associated to $\mathcal{F}^{(r)}$. Then the following hold:
  1. i)
    The total variation measures (see, e.g., [42, §7.2]) $\|T^{(r)}\|$ of $T^{(r)}$ converge to $\mu_{E}$ in the weak topology of measures supported on $E$;
  2. ii)
    If $T\in\mathscr{S}_{0}^{k}(E)$ is an accumulation point of $\{T^{(r)}\}$ in $\mathscr{M}_{0}^{k}(E)$, then there exists an orthonormal basis $\{\tau_{1},\dots,\tau_{n}\}$ of $\mathbb{R}^{n}$ such that
    $T(\omega)=\frac{1}{\binom{n}{k}}{\sum_{|\alpha|=k}}^{\prime}\int\langle\omega(x);\tau_{\alpha_{1}}\wedge\dots\wedge\tau_{\alpha_{k}}\rangle\mathrm{d}\mu_{E}.$
Note that, by minor modificartions of [22, Cor. 5.1], one can prove the converse of ii) of the above conjecture. In particular, for any orthonormal basis $\{\tau_{1},\dots,\tau_{n}\}$ of $\mathbb{R}^{n}$, there exists a sequence of Fekete currents $\{\mathcal{F}^{(r)}\}=\{\mathcal{F}^{1,r},\dots,\mathcal{F}^{N(r),r}\}$ such that the sequence $T^{(r)}$ of the associated empirical currents satisfy $T^{(r)}\rightharpoonup\frac{1}{\binom{n}{k}}\sum^{\prime}_{|\alpha|=k}\tau_{\alpha_{1}}\wedge\dots\wedge\tau_{\alpha_{k}}\mu_{E}$.
5.3. Least squares fitting over integral admissible $k$-meshes
Proposition 5.
Let $\{\mathcal{T}^{(r)}\}$ be an admissible integral $k$-mesh of constant $C$ for the compact set $E$. Let $w_{i}^{(r)}>0$ for any $i=1,\dots,M(r):=\operatorname*{Card}\mathcal{T}^{(r)}$, $r\in\mathbb{N}$, and denote by $P^{(r)}$ the fitting operator defined in Eq. (10) of Section 2. Then one has
$\|P^{(r)}\|_{\mathrm{op}}\leq C\sqrt{\sum_{i=1}^{M(r)}\frac{w_{i}^{(r)}}{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}.$
Hence the following error estimate holds true for any $\omega\in\mathscr{D}_{0}^{k}(E)$:
(57) $\|\omega-P^{(r)}\omega\|_{0}\leq\left(1+C\sqrt{\sum_{i=1}^{M(r)}\frac{w_{i}^{(r)}}{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}\right)d_{r}(\omega,E),$
where we denoted by $d_{r}(\omega,E):=\inf\{\|\omega-\theta\|_{0},\,\theta\in\mathscr{P}_{r}\Lambda^{k}\}$ the error of best approximation in $\mathscr{P}_{r}\Lambda^{k}$ of $\omega$.
Proof.
Recall that Proposition 3 states that
$\|P\|_{\mathrm{op}}\leq\mathcal{M}:=\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M(r)}\left|\sum_{h=1}^{N}T_{i}(\eta_{h})T(\eta_{h})\right|w_{i}^{(r)}\,,$
where $N:=\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}$. Using the sampling property of admissible integral $k$-meshes, we derive
$\mathcal{M}=$ $\sup_{T\in\mathscr{A}^{k}(E)}\sum_{i=1}^{M(r)}\left|T\sum_{h=1}^{N}T_{i}(\eta_{h})\eta_{h}\right|w_{i}^{(r)}\leq C\sum_{i=1}^{M(r)}\max_{j\in\{1,\dots,M(r)\}}\left|T_{j}\sum_{h=1}^{N}T_{i}(\eta_{h})\eta_{h}\right|w_{i}^{(r)}$
$\leq$ $\sum_{i=1}^{M(r)}\max_{j\in\{1,\dots,M(r)\}}\left(\sum_{k=1}^{N}T_{j}^{2}(\eta_{k})\right)^{1/2}\left(\sum_{h=1}^{N}T_{i}^{2}(\eta_{h})\right)^{1/2}w_{i}^{(r)}$
$=$ $C\left(\max_{j\in\{1,\dots,M(r)\}}\sum_{k=1}^{N}T_{j}^{2}(\eta_{k})\right)^{1/2}\sum_{i=1}^{M(r)}\left(\sum_{h=1}^{N}T_{i}^{2}(\eta_{h})\right)^{1/2}w_{i}^{(r)}$
$\leq$ $\frac{C}{\sqrt{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}\left(\sum_{j}^{M(r)}\sum_{k=1}^{N}T_{j}^{2}(\eta_{k})w_{j}^{(r)}\right)^{1/2}\sum_{i=1}^{M(r)}\left(\sum_{h=1}^{N}T_{i}^{2}(\eta_{h})\right)^{1/2}w_{i}^{(r)}$
$\leq$ $\frac{C}{\sqrt{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}\left(\sum_{j}^{M(r)}\sum_{k=1}^{N}T_{j}^{2}(\eta_{k})w_{j}^{(r)}\right)^{1/2}\left(\sum_{i=1}^{M(r)}\sum_{h=1}^{N}T_{i}^{2}(\eta_{h})w_{i}\right)^{1/2}\sqrt{\sum_{i=1}^{M(r)}w_{i}^{(r)}}$
$=$ $C\frac{\sqrt{\sum_{i=1}^{M(r)}w_{i}^{(r)}}}{\sqrt{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}\left(\sum_{k=1}^{N}\sum_{j}^{M(r)}T_{j}^{2}(\eta_{k})w_{j}^{(r)}\right)$
$=$ $C\sqrt{\sum_{i=1}^{M(r)}\frac{w_{i}^{(r)}}{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}N.$
The claim is proved. ∎
Note that, in the case of equal weights $w_{i}^{(r)}\equiv 1/M^{(r)}$, inequality (57) simplifies to
$\|\omega-P^{(r)}\omega\|_{0}\leq\left(1+C\sqrt{M(r)}\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k})\right)d_{r}(\omega,E).$
The error estimate of Proposition 5 is satisfactory for most applications, but quite pessimistic. Indeed, following the lines of [24, Thm. 2], it is possible to obtain a much sharper result:
Corollary 4.
In the setting of Proposition 5, one has
$\|P^{(r)}\omega\|_{0}\leq C\left(\|\omega\|_{0}+\sqrt{\sum_{i=1}^{M(r)}\frac{w_{i}^{(r)}}{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}d_{r}(\omega,E)\right),\quad\forall\omega\in\mathscr{D}_{0}^{k}(E),$
and
(58) $\|\omega-P^{(r)}\omega\|_{0}\leq\left(1+C\left(1+\sqrt{\sum_{i=1}^{M(r)}\frac{w_{i}^{(r)}}{\min_{i=1,\dots,M(r)}w_{i}^{(r)}}}\right)\right)d_{r}(\omega,E),\quad\forall\omega\in\mathscr{D}_{0}^{k}(E)\,.$
From (58) one deduces the advantage of using admissible integral $k$-meshes in the design of an approximation scheme. Considering normalized equal weights, i.e. $w_{i}^{(r)}\equiv 1/M(r)$, one indeed has
$\|\omega-P^{(r)}\omega\|_{0}\leq\left(1+C\left(1+\sqrt{M(r)}\right)\right)d_{r}(\omega,E)\sim 2\widetilde{C}\sqrt{M(r)}d_{r}(\omega,E),\quad\forall\omega\in\mathscr{D}_{0}^{k}(E).$
This last estimate, for a quasi optimal admissible integral $k$-mesh (i.e., $M(r)\sim\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}\log r$) as the one constructed on the simplex by the Algorithm 1 of Section 4, implies
$\|\omega-P^{(r)}\omega\|_{0}\sim 2\widetilde{C}\sqrt{\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}\log r\,}d_{r}(\omega,E),$
while, for the case of an optimal admissible integral $k$-mesh (i.e., $M(r)\sim\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}\log r$) as that of Example 8, the estimate (58) reads
$\|\omega-P^{(r)}\omega\|_{0}\sim 2\widetilde{C}\sqrt{\operatorname*{dim}\mathscr{P}_{r}\Lambda^{k}\,}d_{r}(\omega,E)\,,$
see Remark 14. These error estimates improve the Lebesgue bound (56) for Fekete currents, where the dimension of $\mathscr{P}_{r}\Lambda^{k}$ enters linearly.
Appendix A Integration of differential forms
In the language of differential geometry, an orientation of the real $k$-dimensional manifold $S$ is an equivalence class of oriented atlases of $S$. We recall that two overlapping charts of the manifold $S$ are equioriented if the determinant of the Jacobian of the change of coordinates is positive, and an atlas of $S$ is termed oriented if any pair of overlapping charts in the atlas is equioriented. Existence of an oriented atlas for $S$ is a non trivial fact: if $S$ admits an orientation, then it is said an orientable manifold and the possible orientations of $S$ are the positive (i.e., the one of the oriented atlas we are implicitly considering) and the negative one. Accordingly, a basis $\tau_{1},\dots,\tau_{k}$ of the tangent space $T_{x}S$ of $S$ at $x$ is termed positively oriented or not, depending on the sign of the determinant of the change of basis from $\partial_{1}|_{x},\dots,\partial_{k}|_{x}$ to $\tau_{1},\dots,\tau_{k}$.
A $\mathscr{C}^{m}$ $k$-differential form on real $k$-dimensional smooth manifold $S$ is a $\mathscr{C}^{m}$-smooth section of the $k$-th exterior power of the cotangent bundle of $S$. If $S$ is oriented by the atlas $\{(U_{\alpha},\phi_{\alpha})\}$, it is possible to define the integral of $\omega$ on the domain of a chart $\phi_{\alpha}:S\supset U_{\alpha}\rightarrow V_{\alpha}\subset\mathbb{R}^{k}$ by setting
$\int_{U_{\alpha}}\omega:=\int_{\phi_{\alpha}(U_{\alpha})}(\phi^{-1})^{*}\omega=\int_{V_{\alpha}}\omega(\phi_{\alpha}^{-1}(y))(\mathrm{d}\phi_{\alpha}^{-1}(y)e_{1},\dots,\mathrm{d}\phi_{\alpha}^{-1}(y)e_{k})\mathrm{d}y_{1}\dots\mathrm{d}y_{k}\,,$
being the latter term simply the integral of a smooth function on an open set of $\mathbb{R}^{k}.$ Then, the integral over $S$ of a compactly supported $k$-form is obtained by working with a suitable partition of unity $\{\rho_{\alpha}\}$ subordinated to the considered atlas:
$\int_{S}\omega:=\sum_{\alpha}\int_{U_{\alpha}}\rho_{\alpha}\omega.$
If $S$ is a smooth oriented embedded submanifold of $\mathbb{R}^{n}$, and $\iota:S\hookrightarrow\mathbb{R}^{n}$ denotes the inclusion map, it is customary to speak about integration over $S$ of a $k$-form $\omega$ on $\mathbb{R}^{n}$ instead of the integral of its pullback $\iota^{*}\omega$ with respect to the inclusion map:
$\int_{S}\omega:=$ $\int_{S}\iota^{*}\omega=\sum_{\alpha}\int_{U_{\alpha}}\rho_{\alpha}\iota^{*}\omega=\sum_{\alpha}\int_{V_{\alpha}}(\phi^{-1})^{*}\iota^{*}\omega=\sum_{\alpha}\int_{V_{\alpha}}(\phi^{-1}\circ\iota)^{*}\omega$
$=$ $\sum_{\alpha}\int_{V_{\alpha}}\omega(\tilde{\phi}_{\alpha}^{-1}(y))(\mathrm{d}\tilde{\phi}_{\alpha}^{-1}(y)e_{1},\dots,\mathrm{d}\tilde{\phi}_{\alpha}^{-1}(y)e_{k})\mathrm{d}y_{1}\dots\mathrm{d}y_{k}$
(59) $=$ $\sum_{\alpha}\int_{V_{\alpha}}\left\langle\omega(\tilde{\phi}_{\alpha}^{-1}(y));\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}\right\rangle\mathrm{d}y_{1}\dots\mathrm{d}y_{k},$
where we denoted by $\tilde{\phi}^{-1}$ the map $V_{\alpha}\ni y\mapsto\iota\phi^{-1}(y)\in\mathbb{R}^{n}.$
In the context of geometric measure theory one aims at working with much less regular objects rather than smooth manifolds, hence there is the need for a generalization of Eq. (59). Indeed, one first writes
$\sum_{\alpha}\int_{V_{\alpha}}\left\langle\omega(\tilde{\phi}_{\alpha}^{-1}(y));\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}\right\rangle\mathrm{d}y_{1}\dots\mathrm{d}y_{k},$
$=$ $\sum_{\alpha}\int_{V_{\alpha}}\left\langle\omega(\tilde{\phi}_{\alpha}^{-1}(y));\frac{\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}}{\left|\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}\right|}\right\rangle\left|\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}\right|\mathrm{d}y_{1}\dots\mathrm{d}y_{k}$
$=$ $\sum_{\alpha}\int_{V_{\alpha}}\left\langle\omega(\tilde{\phi}_{\alpha}^{-1}(y));\frac{\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}}{\left|\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}\right|}\right\rangle\llbracket D\tilde{\phi}^{-1}(y)\rrbracket\mathrm{d}y_{1}\dots\mathrm{d}y_{k}\,,$
where $\llbracket A\rrbracket:=\sqrt{\det(A^{\top}A)}$ denotes the Jacobian of the matrix $A$ and the equality $\llbracket D\tilde{\phi}^{-1}(y)\rrbracket=|\partial_{1}\tilde{\phi}^{-1}\wedge\dots\wedge\partial_{k}\tilde{\phi}^{-1}|(y)$ follows from the Cauchy-Binet formula.
Then, using the Area Formula (see, e.g., [30, §3.3]) one notices that the integration on $V_{\alpha}$ with respect to the measure $\llbracket D\tilde{\phi}^{-1}(y)\rrbracket\mathrm{d}y_{1}\dots\mathrm{d}y_{k}$ is precisely the integral over $U_{\alpha}$ with respect to the $k$-dimensional Hausdorff measure $\mathcal{H}^{k}$ restricted to $S$:
(60) $\int_{S}\omega=\int_{S}\langle\omega(x);\tau(x)\rangle\mathrm{d}\mathcal{H}^{k}(x)\,,$
where we set
$\tau(x):=\frac{\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}}{\left|\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{1}}\wedge\dots\wedge\frac{\partial\tilde{\phi}^{-1}(y)}{\partial y_{k}}\right|}\,\Bigg{|}_{y=\tilde{\phi}(x)}\,.$
Since the integrand at the right hand side of the above equation does not change value if we replace $\tau(x)$ with any other unit $k$-vector $\eta:=\eta_{1}\wedge\dots\wedge\eta_{k}$ such that $\{\eta_{1}(x),\dots,\eta_{k}(x)\}$ is a positively oriented basis of $T_{x}S$, the definition of orientation of a $k$-submanifold of $\mathbb{R}^{n}$ used in the context of geometric measure theory is the datum of a continuous $k$-vector field $S\ni x\mapsto\eta(x)\in\Lambda_{k}$ such that, for any $x\in S$, $\eta(x)=\eta_{1}(x)\wedge\dots\wedge\eta_{k}(x)$ is a unit simple vector and $T_{x}S=\operatorname*{span}\{\eta_{1}(x),\dots,\eta_{k}(x)\}.$
Equation (60) can be taken as definition of the integral of $\omega$ over $S$. The advantage of this approach when dealing with $k$-dimensional surfaces that are less regular is ready at hand. Assume that $S$ is a $k$-rectifiable set that we can write as the disjoint union $S=S_{\mathrm{reg}}\cup S_{\mathrm{sing}}$ of the $\mathscr{C}^{1}$ orientable submanifold $S_{\mathrm{reg}}$ and the set $S_{\mathrm{sing}}$ of zero $k$-dimensional Hausdorff measure. Then the right hand side of equation (60) still makes sense on the orientable $k$-rectifiable set $S$, being $\tau$ any extension of the orientation of $S_{\mathrm{reg}}$. Hence one sets
$\int_{S}\omega:=\int_{S_{\mathrm{reg}}}\omega.$